### v1

# Modified Lepage-type test statistics for the weak null hypothesis

Abstract
Detecting simultaneous shifts in the location and scale of two populations is a challenging problem in statistical research. A common way to address this issue is by combining location and scale test statistics. A well-known example is the Lepage test, which combines the Wilcoxon-Mann-Whitney test for location with the Ansari-Bradley test for scale. However, the Wilcoxon-Mann-Whitney test assumes that the population variances are equal, while the Ansari-Bradley test assumes the population medians are equal. This study introduces new approaches that combine recent methodological advances to relax these assumptions. We incorporate the Fligner-Policello test, a distribution-free alternative to the Wilcoxon-Mann-Whitney test that does not require the assumption of equal variances. The Fligner-Policello test is further enhanced by the Fong-Huang method, which provides an improved variance estimation. Additionally, we propose a new variance estimator for the Ansari-Bradley test, thereby eliminating the need for the equal medians assumption. These methodological modifications are integrated into the Lepage framework to operate under a weak null hypothesis. Simulation results suggest that these new tests are promising candidates for location-scale testing. The practical utility of the proposed tests is then demonstrated through an analysis of four real-world biomedical datasets. These empirical applications confirm the robustness and reliability of the modified tests for the two-sample independent location-scale problem.
Keywords: Fligner-Policello estimator; Fong-Huang estimator; Lepage-type statistics; Location-scale model; power of test; variance estimators. 11footnotetext: corresponding author: abid0100@gmail.com
1 Introduction
The comparison of treatment and control groups is a common practice in biomedical research and related disciplines. Standard two-sample tests, such as the Student’s tt–test and FF–ratio test, are frequently employed to evaluate the effects of different treatments, such as varying medication doses, on two independent groups. When the populations are normally distributed, these tests are considered the uniformly most powerful unbiased tests. However, the underlying assumption of normality for these tests is often not met in practice. Consequently, nonparametric two-sample tests are recommended, particularly in situations involving small sample sizes, where the verification of distributional assumptions is challenging or infeasible. These tests are preferred due to their lack of distributional assumptions within each sample and their demonstrated robustness under diverse conditions, as discussed by Blair and Higgins (1980) and Van der Vaart (2000). Among the numerous available statistical tests for comparing the locations of two independent samples, the nonparametric Wilcoxon-Mann-Whitney UU–test is frequently applied in a variety of settings (Letshedi et al. 2021; Lin et al. 2021 and Dao 2022). Ansari-Bradley test is a nonparametric test that is commonly used to compare the dispersion of two independent samples. It has become one of the core statistically analytical tools in most research and practical applications (Mukherjee and Marozzi 2019; Murakami and Neuhäuser 2025 and Hussain 2025).
Simultaneous variation in location and scale is a common occurrence in statistical analysis (see e.g., Neuhäuser 2002 and Murakami and Neuhäuser 2025). For instance, in randomized controlled trials, the null hypothesis of no treatment effect often assumes homogeneous variances among the different treatment groups. However, the presence of heteroscedasticity can itself be an indicator of an underlying treatment effect (Neuhäuser, 2002). A compelling example can be found in a clinical trial analyzed by Neuhäuser (2001), which investigated the frequency of exacerbations in patients with chronic obstructive pulmonary disease (COPD). In this context, a treatment effect could manifest a shift in both the location and scale parameters. This observation has spurred significant research into developing distribution-free tests capable of simultaneously detecting differences in both location and scale between two populations. A prominent example of such a test is the sum of the squared standardized Wilcoxon-Mann-Whitney UU–test and Ansari-Bradley CC–test statistics, a method introduced by Lepage (1971).
Statistical testing typically focuses on the strong null hypothesis, which posits an exact equality between populations. However, a weak null hypothesis, which concerns a relative marginal effect, can be relevant in certain situations (Brunner et al. 2002 and Fong and Huang 2019). The Wilcoxon-Mann-Whitney UU–test, for instance, cannot directly test this weak null hypothesis because the sampling distribution of its test statistic is not distribution-free under this condition (Chung and Romano 2013 and Chung and Romano 2016). To address this limitation, Fligner and Policello (1981) developed an alternative to the UU–test. They introduced a data-based approximation for the variance of the ranks. Their resulting test, known as the Fligner-Policello (FP) test, is calculated using the statistic $(U-1/2)/\sqrt{\widehat{Var}_{FP}(U)}$. Here, $\widehat{Var}_{FP}(U)$ represents a consistent variance estimator, which will be described in detail in Section 2. This test statistic can be approximated by a standard normal distribution or by a permutation distribution. When an additional distributional symmetry condition is met for both random variables $X$ and $Y$, the weak null hypothesis becomes equivalent to testing for the equality of population medians. Furthermore, the FP test is generally preferred over the UU–test when testing the weak null hypothesis without the assumption of homogeneous variances (e.g., Siegel and Castellan 1988; Zumbo and Coulombe 1997 and Mickelson 2013). The FP test, therefore, provides a robust alternative for comparing population locations under less restrictive assumptions.
In this paper, we propose several modifications to the Lepage test statistic to improve its power for the two-sample location-scale problem. These changes are specifically designed to maintain or increase statistical power under weak null hypotheses-that is, when unequal variances exist in location-testing or unequal medians exist in scale-testing. One of our approaches introduces the FP test as an alternative to the traditional Wilcoxon-Mann-Whitney for assessing the location component of the Lepage test. Further, we incorporate two key methodological advancements in variance estimation under less restrictive assumptions. First, we utilize the Fong-Huang (FH) estimator, which corrects a known error in the original FP estimator (Fong and Huang, 2019). This FH test enhances the accuracy to detect the location-difference, allowing it to be reliably compared to standard normal approximation as $(U-1/2)/\sqrt{\widehat{Var}_{FH}(U)}$. Second, we propose a novel empirical variance estimator for the Ansari-Bradley $C$ statistic. It can be implemented through standard normal approximation as $\big(C-E_{0}(C)\big)/\sqrt{\widehat{Var}(C)}$, eliminates the need for the assumption of equal medians. The Lepage-type test statistics developed in this study are constructed by incorporating the variance estimators of the $U$ and $C$ statistics. The subsequent sections of this paper provide the specific details and derivations of these new test statistics.
This study is structured as follows: In Section 2, we begin with the details of the $U$ statistic, its variance, and providing an overview of existing estimators proposed by Fligner and Policello (1981) and Fong and Huang (2019). Section 3 introduces the $C$ statistic, its variance, and a newly proposed variance estimator. Section 4 then presents the location-scale model and the Lepage-type statistics that form the basis of our analysis. Section 5 employs a Monte Carlo simulation to evaluate the performance of our proposed tests with classical Lepage test. In Section 6, we demonstrate the practical application of these tests using four real-world datasets. Finally, Section 7 summarizes the key findings of the research along with some future directions.
2 The $U$ statistic
2.1 The classical $U$ statistic
The Wilcoxon-Mann-Whitney UU–test statistic is based on two independent random samples: $X=(X_{1},X_{2},\ldots,X_{m})$ (independent and identically distributed (i.i.d.) observations from a continuous population) and $Y=(Y_{1},Y_{2},\ldots,Y_{n})$ (i.i.d. observations from another continuous population). Let the distribution functions of $X$ and $Y$ be denoted by $\mathbb{F}$ and $\mathbb{G}$, respectively. The $U$ statistic can be computed using either the sum of ranks of the $Y$ observations in the combined sample or a $U$ statistic
$U=\frac{1}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}I\big(X_{i}<Y_{j}\big).$ (1)
A trade-off exists between computational efficiency and clarity of statistical properties when considering the rank sum and $U$ statistic formulations. These two forms are theoretically equivalent, differing only by constant factors (see, for example, Hollander et al. 2014). While the rank sum formulation is computationally advantageous, expressing the statistic as a $U$ statistic, as in Eq. (1), highlights its consistency and minimum variance unbiasedness in estimating $P(Y>X)$ (Lehmann, 1951).
2.2 Variance of $U$ statistic
The $U$ statistic is a nonparametric test, therefore it does not rely on specific distributional assumptions for the independent variables $X$ and $Y$. Under the null hypothesis ($X$ and $Y$ are identically distributed), the distribution of the $U$ statistic is solely determined by the sample sizes $m$ and $n$. Specifically, $E_{0}(U)=1/2$ and the variance of $U$ statistic is
$Var_{0}(U)$ $=$ $\frac{1}{(mn)^{2}}Var\bigg(\sum_{i=1}^{m}\sum_{j=1}^{n}I\big(X_{i}<Y_{j}\big)\bigg)$
$=$ $\frac{1}{(mn)^{2}}\sum_{i=1}^{m}\sum_{j=1}^{n}\sum_{i^{*}=1}^{m}\sum_{j^{*}=1}^{n}Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i^{*}}<Y_{j^{*}}\big)\bigg).$
The indices $i$ and $j$, and $i^{*}$ and $j^{*}$ can relate in four distinct ways. When $i\neq i^{*}$ and $j\neq j^{*}$, the indicator functions $I(X_{i}<Y_{j})$ and $I(X_{i^{*}}<Y_{j^{*}})$ are independent. Consequently, the total variance can be decomposed into three additive components
$Var_{0}(U)$ $=$ $\frac{1}{(mn)^{2}}\sum_{i=1}^{m}\sum_{j\neq j^{*}}^{n}Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i}<Y_{j^{*}}\big)\bigg)$ (2)
$+\frac{1}{(mn)^{2}}\sum_{i\neq i^{*}}^{m}\sum_{j=1}^{n}Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i^{*}}<Y_{j}\big)\bigg)$
$+\frac{1}{(mn)^{2}}\sum_{i=1}^{m}\sum_{j=1}^{n}Var\bigg(I\big(X_{i}<Y_{j}\bigg).$
This is made explicit by noting that
$Cov\bigg(I\big(X_{i}<Y_{j}\big),I\big(X_{i}<Y_{j^{*}}\big)\bigg)$ $=$ $E\bigg(I\big(X_{i}<Y_{j}\big)I\big(X_{i}<Y_{j^{*}}\big)\bigg)-E\bigg(I\big(X_{i}<Y_{j}\big)\bigg)E\bigg(I\big(X_{i}<Y_{j^{*}}\big)\bigg)$
$=$ $E\bigg(E\big(I(X_{i}<Y_{j})I(X_{i}<Y_{j^{*}})|X_{i}\big)\bigg)-E\bigg(E\big(I(X_{i}<Y_{j})|X_{i}\big)\bigg)^{2}$
$=$ $E\bigg(S_{Y}^{2}(X_{i})\bigg)-E^{2}\bigg(S_{Y}(X_{i})\bigg),$
where $S_{Y}(X_{i})=E_{Y|X_{i}}\big(I(X_{i}<Y)\big)$. Thus we have
$Var_{0}(U)$ $=$ $\frac{1}{(mn)^{2}}\bigg(m(n^{2}-n)Var_{0}\big(S_{Y}(X_{i})\big)+n(m^{2}-m)Var_{0}\big(S_{X}(Y_{i})\big)+mnVar_{0}\big(I(X_{i}<Y_{j})\big)\bigg)$ (3)
$=$ $\frac{1}{(mn)^{2}}\bigg(m(n^{2}-n)Var_{0}\big(F_{Y}(X_{i})\big)+n(m^{2}-m)Var_{0}\big(F_{X}(Y_{i})\big)+mnVar_{0}\big(I(X_{i}<Y_{j})\big)\bigg)$
$=$ $\bigg(1-\frac{1}{n}\bigg)\frac{1}{m}Var_{0}\big(F_{Y}(X_{i})\big)+\bigg(1-\frac{1}{m}\bigg)\frac{1}{n}Var_{0}\big(F_{X}(Y_{i})\big)+\frac{1}{mn}Var_{0}\big(I(X_{i}<Y_{j})\big),$
where $S_{X}(Y_{i})=E_{X|Y_{i}}\big(I(Y_{i}<X)\big)$. Under the strong null hypothesis, the cumulative distribution functions $F_{Y}(X_{i})$ and $F_{X}(Y_{i})$ are uniformly distributed random variables. Consequently, the expression evaluates to
$Var_{0}(U)$ $=$ $\bigg(1-\frac{1}{n}\bigg)\frac{1}{12m}+\bigg(1-\frac{1}{m}\bigg)\frac{1}{12n}+\frac{1}{4mn}$
$=$ $\frac{n-1}{12mn}+\frac{m-1}{12mn}+\frac{1}{4mn}$
$=$ $\frac{1}{12}\bigg(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\bigg).$
For sufficiently large $m$ and $n$, the standardized statistic $(U-1/2)/\sqrt{Var_{0}(U)}$ asymptotically follows the standard normal (Hollander et al. 2014), enabling approximate pp–value calculations.
2.3 Fligner and Policello estimator
Several procedures exist for testing the equality of two medians when the populations exhibit differing shapes. The method proposed by Fligner and Policello (1981) is particularly useful as it requires fewer assumptions regarding the distributional forms of the two populations. Their estimator for the variance of the $U$ statistic can be expressed as (Fong and Huang 2019)
$\widehat{Var}_{FP}(U)$ $=$ $\frac{(m-1)n^{2}}{(mn)^{2}}\widehat{Var}(\hat{F}_{Y}(X_{i}))+\frac{(n-1)m^{2}}{(mn)^{2}}\widehat{Var}(\hat{F}_{X}(Y_{i}))+\frac{mn}{(mn)^{2}}\widehat{Var}\big(I(X_{i}<Y_{j})\big)$
$=$ $\bigg(1-\frac{1}{m}\bigg)\frac{1}{m}\widehat{Var}(\hat{F}_{Y}(X_{i}))+\bigg(1-\frac{1}{n}\bigg)\frac{1}{n}\widehat{Var}(\hat{F}_{X}(Y_{i}))+\frac{1}{mn}P_{m}\big(\hat{F}_{Y}(X_{i})\big)P_{n}\big(\hat{F}_{X}(Y_{i})\big).$
In the above expression $\widehat{Var}$ is denoted as the estimator of $Var_{0}$. Furthermore, $\widehat{Var}\big(I(X_{i}<Y_{j})\big)$ is given by $P_{m}\big(\hat{F}_{Y}(X_{i})\big)P_{n}\big(\hat{F}_{X}(Y_{i})\big)$, where $P_{m}$ represents the empirical average operator with respect to the $m$ observations $X_{1},X_{2},\cdots,X_{m}$, and $P_{n}$ represents the empirical average operator with respect to the $n$ observations $Y_{1},Y_{2},\cdots,Y_{n}$. In the case where $m\neq n$, this estimator yields the following estimate under the strong null hypothesis.
$\widehat{Var}_{FP}(U)$ $=$ $\big(1-\frac{1}{m}\big)\frac{1}{12m}+\big(1-\frac{1}{n}\big)\frac{1}{12n}+\frac{1}{4mn}$
$=$ $\frac{1}{12}\bigg(\frac{1}{m}+\frac{1}{n}-\frac{1}{m^{2}}-\frac{1}{n^{2}}+\frac{3}{mn}\bigg),$
which is not equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$. However, this difference is negligible, being of order $O\big(m^{-1}n^{-1}\big)$.
2.4 Fong and Huang estimator
Fong and Huang (2019) recently proposed a corrected estimator for the variance of the $U$ statistic (3), defined as
$\widehat{Var}_{FH}(U)$ $=$ $\bigg(1-\frac{1}{n}\bigg)\frac{1}{m}\widehat{Var}\big(\hat{F}_{Y}(X_{i})\big)+\bigg(1-\frac{1}{m}\bigg)\frac{1}{n}\widehat{Var}(\hat{F}_{X}(Y_{i}))+\frac{1}{mn}\widehat{Var}\big(I(X_{i}<Y_{j})\big).$
Assuming the strong null hypothesis, the following estimate is obtained
$\widehat{Var}_{FH}(U)$ $=$ $\big(1-\frac{1}{n}\big)\frac{1}{12m}+\big(1-\frac{1}{m}\big)\frac{1}{12n}+\frac{1}{4mn}$
$=$ $\frac{1}{12}\bigg(\frac{1}{m}-\frac{1}{mn}+\frac{1}{n}-\frac{1}{mn}+\frac{3}{mn}\bigg),$
which is equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$, when $m\neq n$.
3 The $C$ statistic
3.1 The classical $C$ statistic
The $C$ statistic, proposed by Ansari and Bradley (1960), is used to assess a two-sample dispersion problem. Its application requires the two populations to have identical medians. To calculate the statistic, the combined sample of $X$ and $Y$ is first ordered from least to greatest, represented as $\{Z_{1},~Z_{2},~\cdots,~Z_{N}\}$. Ranks are then assigned symmetrically from the tails of the ordered sample. Rank 1 is assigned to the minimum ($Z_{1}$) and maximum ($Z_{N}$) values, rank 2 to the second smallest ($Z_{2}$) and second largest ($Z_{N-1}$), and so on. The array of assigned ranks, denoted as $S_{N}$, depends on whether the combined sample size $N$ is even or odd. If $N$ is an even integer, the ranks are given by
$S_{N}=\bigg\{1,~2,~3,~\cdots,~\frac{N}{2},~\frac{N}{2},~\cdots,~3,~2,~1\bigg\}.$
If $N$ is an odd integer, the ranks are given by
$S_{N}=\bigg\{1,~2,~3,~\cdots,\frac{N-1}{2},~\frac{N+1}{2},~\frac{N-1}{2},~\cdots,~3,~2,~1\bigg\}.$
Let $R_{i}$ be the rank assigned to the $i$th value from the $Y$ sample. The test statistic, $C$, is defined as the sum of the ranks assigned to all values in the $Y$ sample
$C=\sum_{i=1}^{n}R_{i},$ (4)
where $n$ is the size of the $Y$ sample. An alternative formulation, equivalent to Eq. (4) and more useful for mathematical considerations, is given by
$C=\sum_{i=1}^{k}i\lambda_{i}+\sum_{i=k+1}^{N}(N+1-i)\lambda_{i},$
where
$\lambda_{i}=\begin{cases}1;&\ \ \text{if $Z_{i}\in$ $Y$ }\\ &\\ 0;&\ \ \text{otherwise,}\end{cases}$
and $k$ is the largest integer in $\frac{N+1}{2}$. Small values of $C$ suggest a larger dispersion for the $Y$ sample, whereas large values of $C$ indicate a larger dispersion for the $X$ sample.
3.2 Variance of $C$ statistic
Under the null hypothesis ($H_{0}$), all $\binom{N}{n}$ possible assignments of ranks to the combined sample are equally probable. For large sample approximation, the statistic $\frac{C}{n}$ represents the average of the ranks assigned to the $Y$ sample. Its distribution is identical to the distribution of the sample mean for a random sample of size $n$ drawn without replacement from the finite population of ranks. For an even $N$, the expected value of the statistic, $C$, is given by
$E_{0}\bigg(\frac{C}{n}\bigg)$ $=$ $\mu=\frac{2}{N}\sum_{i=1}^{N/2}i=\frac{\frac{2}{N}\times\frac{N}{2}\bigg(\frac{N}{2}+1\bigg)}{2}=\frac{N+2}{4}$
$E_{0}(C)$ $=$ $\frac{n(N+2)}{4}.$
The variance of $C$ can be calculated as
$Var_{0}\bigg(\frac{C}{n}\bigg)$ $=$ $\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}Cov(R_{i},R_{j})$
$=$ $\frac{1}{n^{2}}\sum_{i=1}^{n}Var(R_{i})+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j\neq i}Cov(R_{i},R_{j})$
$=$ $\frac{\sigma^{2}}{n}+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j\neq i}\bigg(\frac{-\sigma^{2}}{N-1}\bigg)=\frac{\sigma^{2}}{n}-\frac{n(n-1)}{n^{2}}\bigg(\frac{\sigma^{2}}{N-1}\bigg)$
$Var_{0}(C)$ $=$ $\sigma^{2}\bigg(\frac{n(N-n)}{N-1}\bigg),$ (5)
where
$\sigma^{2}$ $=$ $\frac{2}{N}\sum_{i=1}^{N/2}i^{2}-\mu^{2}=\frac{N^{2}-4}{48}.$
Therefore, the final expression for the variance is
$Var_{0}(C)$ $=$ $\bigg(\frac{N^{2}-4}{48}\bigg)\bigg(\frac{n(N-n)}{N-1}\bigg)=\frac{mn(N^{2}-4)}{48(N-1)}.$ (6)
For an odd $N$, the expected value and variance can easily be calculated as
$E_{0}(C)$ $=$ $\frac{n(N+1)^{2}}{4N}$
$Var_{0}(C)$ $=$ $\frac{mn(N+1)\big(N^{2}+3\big)}{48N^{2}}.$ (7)
These results are consistent with standard theory for the mean of a sample from a finite population (see, for example, Hollander et al. 2014).
For any NN (even or odd), the standardized statistic, $C^{*}$, is asymptotically normally distributed under $H_{0}$
$C^{*}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
Asymptotic normality results for $C^{*}$ are also available under general alternatives to $H_{0}$ (see, for example, Ansari and Bradley 1960 and Randles and Wolfe 1979).
3.3 A suggested variance estimator
This section presents a procedure for testing the dispersion of two independent samples, particularly when the underlying assumption regarding populations medians is violate. A key component of this method involves estimating the population variance, $\sigma^{2}$, using results from a sample. The expected value of this estimator is given by
$E(\hat{\sigma}^{2})=\sigma^{2}\bigg(1-\frac{1}{n}\times\frac{N-n}{N-1}\bigg)=\sigma^{2}\bigg(\frac{N(n-1)}{n(N-1)}\bigg).$
Using the known expression for the variance of $C$ (3.2) and the estimator $\hat{\sigma}^{2}$, the variance of $C$ can be estimated as follows
$\widehat{Var}(C)=\hat{\sigma}^{2}\times\bigg(\frac{n(N-1)}{N(n-1)}\bigg)\times\bigg(\frac{n(N-n)}{N-1}\bigg)=\hat{\sigma}^{2}\times\bigg(\frac{n^{2}(N-n)}{N(n-1)}\bigg),$ (8)
where $\hat{\sigma}^{2}$ represents the observed variance of the ranks of sample $Y$. The proposed estimator is equally applicable for two different variances, as described in Eq. (6) and Eq. (3.2), regardless of whether the sample size $N$ is even or odd.
Following the strategy from Fong and Huang (2019), a simulation study is conducted to evaluate the performance of the proposed variance estimator. Data are generated from two different distributions
  1. $\bullet$
    A logistic distribution: $X\sim\mathcal{L}(0,1)$ versus $Y\sim\mathcal{L}(0,1)$,
  2. $\bullet$
    An exponential distribution: $X\sim\mathcal{E}(1)$ versus $Y\sim\mathcal{E}(1)$.
To compute the true variance of the $C$ statistic, Eq. (6) or (3.2) is used, depending on whether the sample size ($N$) is even or odd. This true variance is then compared to the average of the variance estimates obtained using Eq. (8). Table 1, based on 10,000 replications, shows that the proposed estimator provides accurate and consistent estimates of the true variance for both probability distributions. This consistency suggests that the proposed estimator is a consistent and distribution-free variance estimator of $C$ statistic. The test statistic for the weak null hypothesis is therefore given by
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}.$
Table 1: Comparison of $C$ statistic variance and its estimator.
                                                                  Sample sizes
$m=n=5$ $m=n=10$ $m=20,n=15$ $m=30,n=25$ $m=n=30$ $m=n=50$
$Var_{0}(C)$ 5.5556 43.4211 224.4485 874.1319 1142.7970 5258.8380
Logistic distribution
$\widehat{Var}(C)$ 5.5671 43.4862 224.8931 874.9312 1142.5058 5257.0210
Exponential distribution
$\widehat{Var}(C)$ 5.5507 43.5181 225.4013 873.8417 1143.1610 5256.8125
Using a permutation-based approach, we calculated the exact critical values for the $C^{*}$ and $C_{P}^{*}$ test statistics. Our results indicated that both statistics have identical critical values across various sample size combinations. Specifically, the critical values are found at 5% upper-tailed to be 1.697056 for $m=n=5$, 1.669331 for $m=n=10$, 1.674574 for $m=20,n=10$, and 1.62453 for $m=n=20$. The observed similarity in critical values suggests that the asymptotic normality of the $C_{P}^{*}$ test statistic under the null hypothesis is consistent with that of the $C^{*}$ test statistic.
3.4 Foundation in Slutsky’s theorem
Slutsky’s theorem is a foundational principle in probability theory that is instrumental in establishing the asymptotic normality of a statistic, particularly when it incorporates a consistent estimator. The theorem states that if a sequence of random variables, $A_{n}$, converges in distribution, $D$, and another sequence of random variables, $B_{n}$, converges in probability to a constant, $a$, then the product of these sequences, $A_{n}B_{n}$, converges in distribution to $aD$ (see e.g., Casella and Berger 2002, p. 239–240). This can be expressed as
$\text{If}~~A_{n}\overset{d}{\longrightarrow}D~~\text{and}~~B_{n}\overset{p}{\longrightarrow}a,~~\text{then}~~A_{n}B_{n}\overset{d}{\longrightarrow}aD.$
We can apply this theorem to demonstrate the asymptotic normality of the test statistic, $C_{P}^{*}$. This statistic can be expressed as the product of two components in the following form
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\times\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}.$ (9)
The first component of Eq. (9), $\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}$, converges in distribution to a standard normal distribution, whereas the second component, $\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}$, converges in probability to 1. According to Slutsky’s theorem, their product, the test statistic $C_{P}^{*}$ is asymptotically normally distributed under a weak $H_{0}$
$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
4 Location-scale models
4.1 Hypotheses
Let $X=(X_{1},X_{2},\ldots,X_{m})$ and $Y=(Y_{1},Y_{2},\ldots,Y_{n})$ denote two i.i.d. random samples of sizes $m$ and $n$, respectively. Let $\mathbb{F}$ and $\mathbb{G}$ represent the continuous distribution functions associated with the samples $X$ and $Y$, respectively. The null hypothesis under consideration posits that both samples are drawn from the same population as
$H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta),~~~\forall~\theta\in(-\infty,\infty).$
Our aim is to assess the null hypothesis against the location-scale alternative
$H_{1}:\mathbb{F}(\theta)=\mathbb{G}\left(\frac{\theta+\Delta}{\tau}\right),$
where
$\Delta\neq 0,~~~\tau\neq 1,~~~\Delta\in(-\infty,\infty),~~~\tau\in(0,\infty).$
In hypothesis $H_{1}$, $\Delta$ represents the location difference, while $\tau$ represents the scale difference between the distributions $\mathbb{F}$ and $\mathbb{G}$.
4.2 The Lepage-type test statistics
The classical Lepage statistic, introduced by Lepage (1971), combines the Wilcoxon-Mann-Whitney UU–test for location difference and the Ansari-Bradley CC–test for scale variation. The $U$ statistic, defined in Eq. (1), is applicable to the location model $\mathbb{F}(\theta)=\mathbb{G}(\theta+\Delta)$, as formalized by the following hypotheses
$H_{0}:\Delta=0~~~~~~\text{versus}~~~~~~H_{1}:\Delta<0,~\Delta>0~~\text{or}~~\Delta\neq 0.$
The CC–test, within the framework of a location-scale model, relates to the scale model $\mathbb{F}(\theta)=\mathbb{G}(\theta/\tau)$ by imposing the restriction that the location parameter $\Delta$ equals zero. This restriction gives the subsequent hypotheses
$H_{0}:\tau=1~~~~~~\text{versus}~~~~~~H_{1}:\tau<1,~\tau>1~~\text{or}~~\tau\neq 1.$
Then Lepage statistic is defined as
$L_{0}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}.$
Lepage demonstrated that the statistics $U$ and $C$ are linearly uncorrelated inside the $L_{0}$ statistic framework for all $m$ and $n$; hence, this test under $H_{0}$ asymptotically conforms to a $\chi^{2}$ distribution with 2 degrees of freedom (Lepage, 1971).
This research presents a methodology for constructing five Lepage-type test statistics to assess a weak null hypothesis. These statistics are constructed using the two estimators for $Var_{0}(U)$ detailed in Section 2 and the single estimator for $Var_{0}(C)$ described in Section 3. These estimators are formulated within the framework of the Lepage test statistic as follows
$L_{1}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
$L_{2}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
$L_{3}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
$L_{4}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
and
$L_{5}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}.$
Under the null hypothesis ($H_{0}$), the test statistic $L_{i}$ (for all $i=1,2,\cdots,5$) is assumed to follow an asymptotic chi-squared distribution with two degrees of freedom. This is because the constituent statistics, $U$ and $C$, are independent of each other and each of them asymptotically follows a standard normal distribution, as discussed in Sections 2 and 3.
4.3 Critical values for small sample sizes
The Lepage-type test statistics, when used with a normal approximation, are expected to maintain the nominal Type I error rate under the null hypothesis, provided that the sample sizes are sufficiently large. However, for small sample sizes, a permutation-based methodology offers a more robust approach for controlling the Type I error rate. To construct the permutation distribution, an ideal approach would be to enumerate all possible ways of reassigning group labels among the $N$ total samples and calculate the test statistic for each permutation. Due to the computational infeasibility of exhaustive enumeration for even moderately sized samples, a Monte Carlo approach is typically employed. This method involves generating a large number of random permutations to approximate the null distribution.
In this study, we utilized the R package NSM3, which is specifically designed to provide exact and Monte Carlo permutation-based methods for various nonparametric statistical tests, including the Lepage test. The function $cLepage(\alpha,m,n)$” is used to compute the critical values. For all modified Lepage test statistics, we conducted a simulation study under small-sample scenario, with 100,000 simulations. To ensure the consistency of the results, we varied the underlying probability distributions under the null hypothesis, including the normal, uniform, logistic, and Laplace distributions. Across all distributions, the resulting critical values remained consistent. These values are presented in Table 2.
Table 2: Critical values for right-tailed Lepage-type test statistics at a 5% significance level.
                                                       Test statistic
$m$ $n$ $L_{0}$ $L_{1}$ $L_{2}$ $L_{3}$ $L_{4}$ $L_{5}$
5 5 5.3345 8.8948 8.8948 7.2012 10.3906 10.3906
6 5 5.5269 7.7793 7.9803 7.7727 12.4460 12.4467
6 6 5.7692 6.8571 6.8571 6.8173 11.2084 11.2084
7 5 5.5720 7.9068 7.9068 8.5803 11.5886 11.5886
7 7 5.6541 6.8855 6.8855 7.0367 9.0802 9.0802
8 5 5.5037 7.5440 7.4924 8.6301 12.2734 12.2084
8 8 5.6775 6.6280 6.6280 6.8773 8.3711 8.3711
9 5 5.4444 7.2574 7.2574 8.3956 12.0756 12.0756
10 5 5.4468 7.3250 7.3355 9.0133 11.6773 11.6655
10 10 5.7436 6.5368 6.5719 6.5545 7.5905 7.5905
5 Size and power comparison
A Monte Carlo simulation study is conducted in R (version 4.5.1) to assess the performance of Lepage-type test statistics. Power levels are estimated based on 10,000 simulations, with the significance level for rejecting the null hypothesis ($H_{0}$) set at $0.05$. The power performances of the proposed $L_{i}$ (for all $i=1,2,\cdots,5$) test statistics are compared to that of the classical $L_{0}$ test under simultaneous location-scale shift. For the sake of brevity, the significance level and estimated power of $L_{0}$ and $L_{i}$ (for all $i=1,2,\cdots,5$) are examined for two sample size combinations, $(m,n)$: $(10,10)$ and $(40,30)$. The hypotheses under consideration are $H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta)$ versus $H_{1}:\mathbb{F}(\theta)=\mathbb{G}(\frac{\theta-\Delta}{\tau})$, where $\Delta\neq 0$ and $\tau\neq 1$. To evaluate the performance of the competing Lepage-type test statistics, $\mathbb{F}$ and $\mathbb{G}$ are used to represent the following distributions.
  1. $\bullet$
    $\mathbb{F}\sim\mathcal{E}(0.5)$ versus $\mathbb{G}\sim\mathcal{E}(\theta)$: The exponential distribution is a continuous probability distribution characterized by its right-skewed nature. It is frequently employed to model the duration of time until a specific event occurs, such as the time until a mechanical component experiences failure, the inter-arrival time of customers at a service point, or the decay time of a radioactive particle.
  2. $\bullet$
    $\mathbb{F}\sim\chi^{2}_{(1)}$ versus $\mathbb{G}\sim\chi^{2}_{(\theta)}$: The chi-square distribution is a continuous, asymmetric probability distribution characterized by a rightward skew. It is a fundamental tool in inferential statistics, particularly for applications in hypothesis testing.
  3. $\bullet$
    $\mathbb{F}\sim\mathcal{G}(1,1)$ versus $\mathbb{G}\sim\mathcal{G}(1,\theta)$: The Gamma distribution is a versatile, continuous probability distribution characterized by a rightward skew. It is widely employed to model waiting times and other non-negative random variables, finding extensive application in diverse fields, such as reliability engineering, queueing theory, insurance and hydrology, etc.
  4. $\bullet$
    $\mathbb{F}\sim\mathcal{B}(2,2)$ versus $\mathbb{G}\sim\mathcal{B}(2,\theta)$: The beta distribution is a family of continuous probability distributions defined on the interval $[0,1]$. It is primarily utilized to model the behavior of random variables representing proportions, percentages, or probabilities. Unlike unimodal distributions like the normal or Laplace distributions, the beta distribution exhibits remarkable flexibility in its shape. Its form can vary significantly, ranging from a straight line to U-shaped, J-shaped, or even bell-shaped, depending on the specific values of its two parameters. This adaptability makes it a valuable tool in various fields. The beta distribution is well-suited for several applications, including Bayesian inference, project management and reliability engineering.
  5. $\bullet$
    $\mathbb{F}\sim\mathcal{U}(0,1)$ versus $\mathbb{G}\sim\mathcal{U}(0,\theta)$: A continuous uniform distribution is a probability distribution where every outcome within a defined interval has an equal probability of occurrence. This core property makes it a valuable baseline model in statistical analysis and simulations. The uniform distribution is widely used in situations where there is no a priori reason to believe that any single outcome is more likely than another. It is widely employed for statistical analysis, simulations, and the modeling of real-world phenomena.
Simulation results for the exponential distribution are presented in Table 3, examining both small $(m=n=10)$ and large $(m=40,n=30)$ sample sizes. The first column, which corresponds to the case of identical distributions, shows the simulated Type I error rate. A size correct test must have a Type I error rate close to the nominal significance level of $0.05$. For large samples, the $L_{0}$ test appears slightly conservative, while the proposed tests are slightly liberal. This is likely due to the use of asymptotic critical values rather than permutation-based ones. All other columns demonstrate that the proposed tests have significantly higher power than $L_{0}$, with the exception of the $L_{3}$ test. These results suggest that the proposed Lepage-type tests are a suitable alternative for analyzing exponentially distributed data.
Simulation results for the chi-square distribution are displayed in Table 4. For small samples, the Type I error rates are close to the nominal level. However, for large samples, a slightly liberal behavior is observed in the $L_{3}$, $L_{4}$ and $L_{5}$ test statistics. The power comparison shows that the proposed tests consistently maintain higher power across all findings for both small and large sample sizes. As observed previously with the exponential distribution, the proposed $L_{3}$ test statistic demonstrates the lowest power.
An analysis of the Gamma distribution is presented in Table 5. The simulation study indicates that the proposed statistics consistently outperform the classic $L_{0}$ statistic for both small and large sample comparisons, with the exception of $L_{3}$. For Type I error, all tests maintain the rate near $0.05$, though $L_{4}$ and $L_{5}$ exhibit a slightly liberal behavior in the large sample comparison. In summary, the proposed test statistics generally demonstrate greater efficiency than the $L_{0}$ statistic for estimating parameters of the Gamma distribution under both small and large sample conditions.
Table 3: Simulated actual Type I error rate and power of $\mathcal{E}(0.5)$ versus $\mathcal{E}(\theta)$.
$\theta$
$m$ $n$ Test 0.5 1.0 1.5 2.0 2.5
10 10 $L_{0}$ 0.0526 0.1891 0.4071 0.5992 0.7343
$L_{1}$ 0.0509 0.2169 0.4758 0.6680 0.7829
$L_{2}$ 0.0473 0.2127 0.4665 0.6582 0.7801
$L_{3}$ 0.0544 0.1630 0.3557 0.5233 0.6684
$L_{4}$ 0.0525 0.1958 0.4295 0.6168 0.7467
$L_{5}$ 0.0513 0.1944 0.4283 0.6217 0.7480
40 30 $L_{0}$ 0.0458 0.6164 0.9587 0.9984 0.9999
$L_{1}$ 0.0544 0.6609 0.9704 0.9987 1.0000
$L_{2}$ 0.0553 0.6568 0.9683 0.9982 1.0000
$L_{3}$ 0.0583 0.6317 0.9603 0.9970 0.9998
$L_{4}$ 0.0608 0.6615 0.9699 0.9980 0.9999
$L_{5}$ 0.0612 0.6676 0.9696 0.9975 0.9998
Table 4: Simulated actual Type I error rate and power of $\chi^{2}_{(1)}$ versus $\chi^{2}_{(\theta)}$.
$\theta$
$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
10 10 $L_{0}$ 0.0534 0.1051 0.2626 0.4508 0.6387
$L_{1}$ 0.0492 0.1279 0.3203 0.5284 0.7110
$L_{2}$ 0.0496 0.1373 0.3124 0.5320 0.7165
$L_{3}$ 0.0503 0.1022 0.2135 0.3814 0.5606
$L_{4}$ 0.0536 0.1188 0.2726 0.4802 0.6525
$L_{5}$ 0.0511 0.1146 0.2831 0.4747 0.6642
40 30 $L_{0}$ 0.0480 0.3230 0.7984 0.9738 0.9978
$L_{1}$ 0.0535 0.3582 0.8320 0.9811 0.9986
$L_{2}$ 0.0521 0.3492 0.8265 0.9806 0.9987
$L_{3}$ 0.0584 0.3348 0.7925 0.9772 0.9985
$L_{4}$ 0.0604 0.3709 0.8323 0.9815 0.9987
$L_{5}$ 0.0583 0.3659 0.8309 0.9790 0.9984
Table 5: Simulated actual Type I error rate and power of $\mathcal{G}(1,1)$ versus $\mathcal{G}(1,\theta)$.
$\theta$
$m$ $n$ Test 1.0 2.0 3.0 4.0 5.0
10 10 $L_{0}$ 0.0496 0.1882 0.4070 0.6077 0.7246
$L_{1}$ 0.0490 0.2236 0.4787 0.6627 0.7916
$L_{2}$ 0.0475 0.2220 0.4770 0.6621 0.7812
$L_{3}$ 0.0504 0.1638 0.3523 0.5333 0.6714
$L_{4}$ 0.0533 0.1951 0.4282 0.6200 0.7492
$L_{5}$ 0.0478 0.1962 0.4241 0.6216 0.7501
40 30 $L_{0}$ 0.0496 0.6163 0.9621 0.9971 0.9997
$L_{1}$ 0.0500 0.6651 0.9712 0.9985 1.0000
$L_{2}$ 0.0494 0.6504 0.9677 0.9989 1.0000
$L_{3}$ 0.0536 0.6255 0.9633 0.9975 0.9999
$L_{4}$ 0.0607 0.6713 0.9716 0.9982 0.9999
$L_{5}$ 0.0612 0.6735 0.9731 0.9979 0.9999
In a simulation study of the beta distribution, the proposed test statistics clearly outperform the traditional $L_{0}$ statistic in detecting simultaneous shifts under both small and large sample comparisons, as shown in Table 6. The $L_{4}$ and $L_{5}$ tests are slightly liberal in the large sample comparison. The proposed tests show superior power compared to the classical Lepage test. For this distribution, $L_{3}$ has the lowest power in the small sample investigation but performs similarly to $L_{0}$ in the large sample case. These findings suggest that the proposed tests offer a robust alternative to $L_{0}$ for analyzing parameters of the beta distribution.
For the uniform distribution, all competing tests show similar performance in both sample scenarios, see Table 7. Consistent with previous findings, the proposed $L_{3}$, $L_{4}$, and $L_{5}$ tests are slightly liberal for large sample investigation.
Table 6: Simulated actual Type I error rate and power of $\mathcal{B}(2,2)$ versus $\mathcal{B}(2,\theta)$.
$\theta$
$m$ $n$ Test 2.0 3.0 4.0 5.0 6.0
10 10 $L_{0}$ 0.0495 0.1221 0.2927 0.4647 0.6441
$L_{1}$ 0.0512 0.1414 0.3333 0.5486 0.7104
$L_{2}$ 0.0504 0.1391 0.3413 0.5482 0.7093
$L_{3}$ 0.0487 0.1033 0.2407 0.3971 0.5708
$L_{4}$ 0.0518 0.1252 0.3074 0.4871 0.6692
$L_{5}$ 0.0503 0.1204 0.3059 0.4910 0.6614
40 30 $L_{0}$ 0.0509 0.3688 0.8383 0.9843 0.9989
$L_{1}$ 0.0512 0.4127 0.8698 0.9874 0.9981
$L_{2}$ 0.0511 0.4020 0.8659 0.9852 0.9990
$L_{3}$ 0.0525 0.3737 0.8420 0.9837 0.9988
$L_{4}$ 0.0633 0.4142 0.8763 0.9855 0.9992
$L_{5}$ 0.0577 0.4188 0.8683 0.9856 0.9990
Table 7: Simulated actual Type I error rate and power of $\mathcal{U}(0,1)$ versus $\mathcal{U}(0,\theta)$.
$\theta$
$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
10 10 $L_{0}$ 0.0493 0.2565 0.5619 0.7444 0.8454
$L_{1}$ 0.0491 0.2794 0.5699 0.7462 0.8498
$L_{2}$ 0.0440 0.2813 0.5627 0.7524 0.8545
$L_{3}$ 0.0488 0.2417 0.5213 0.7029 0.8130
$L_{4}$ 0.0480 0.2448 0.5402 0.7318 0.8339
$L_{5}$ 0.0549 0.2556 0.5362 0.7253 0.8359
40 30 $L_{0}$ 0.0435 0.7935 0.9888 0.9993 1.0000
$L_{1}$ 0.0489 0.8122 0.9904 0.9997 1.0000
$L_{2}$ 0.0526 0.8054 0.9912 0.9994 1.0000
$L_{3}$ 0.0572 0.7874 0.9879 0.9994 1.0000
$L_{4}$ 0.0629 0.7997 0.9897 0.9994 0.9999
$L_{5}$ 0.0638 0.7993 0.9894 0.9995 1.0000
6 The empirical study
The focus of this section is to explore the application and evaluation of proposed Lepage-type tests. We compare their performance to the traditional Lepage test using four illustrative datasets, seeking to identify significant changes in location, scale, or both. Figure 1 displays box-plots of the data, and Table 8 provides detailed dataset descriptions.
Our initial dataset, derived from Karpatkin et al. (1981) and subsequently analyzed by Hollander et al. (2014), investigates the effect of maternal steroid treatment (prednisone) on newborn platelet counts. One group of mothers received no prednisone, while the other did. The treatment aims to increase newborn platelet counts, reducing the risk of cerebral hemorrhage, by inhibiting splenic clearance of maternal antibody-coated platelets. We sought to determine if prednisone increases platelet counts. Figure 1 (upper-left panel) suggests increased variability in the treated group, prompting us to use Lepage-type tests to detect simultaneous location-scale shifts. The pp–values reported in Table 9 indicate that all competing tests detected a statistically significant difference at the 5% level of significance. This result is in agreement with the visual observations.
The second dataset is taken from Hollander et al. (2014), also investigated by Marozzi (2009) and Kössler and Mukherjee (2020). This dataset originates from a biomedical experiment designed to compare peak human plasma growth hormone levels following arginine hydrochloride infusion. Group A ($m=10$), characterized by heightened cardiovascular risk and Type A personality traits, is compared with Group B ($n=11$), exhibiting resistance to coronary disease. Visual inspection (Figure 1, upper-right panel) reveals substantial inter-group differences in skewness, tail weight, location, and scale, justifying a location-scale test. The $L_{0}$, and the proposed tests, applied to this data, all yielded pp–values below 0.05 (Table 9), confirming the visual observations and aligning with prior findings (see, for example, Marozzi 2009 and Kössler and Mukherjee 2020).
The effect of thyroxine on juvenile mice growth is explored in Dataset 3 (Tasdan and Sievers, 2009). Visual comparison of the control $(m=7)$ and treatment $(n=8)$ groups’ weight measurements (daily for 30 days) in Figure 1 (bottom-left panel) reveals marked differences in distribution, including skewness, tail weight, location, and scale. These distributional characteristics suggest of a location-scale test. Therefore, $L_{0}$, and the proposed tests are employed to analyze this weight data. The resulting pp–values (Table 9) indicate statistically significant differences in weight between the control and treatment groups at the 5% significance level for all three tests.
We investigated the sensitivity of location-scale tests to detect subtle differences in cholesterol levels related to smoking status, using Dataset 4 from the Data Story Library, a dataset also analyzed by Kössler and Mukherjee (2020). This dataset compares cholesterol levels in age-matched ex-smokers (cessation within five years) and long-term smokers (25+ years). Our research question focused on whether the Lepage-type tests could identify minor variations, even when no significant differences in mean or variance are observed, as visualized in Figure 1 (bottom-right panel). As shown in Table 9, the insignificant pp–values obtained from all competing tests indicate no difference in cholesterol levels between the groups, demonstrating the effectiveness of our proposed approaches and their alignment with traditional method.
Table 8: Real life datasets.
Dataset 1: Platelet counts of newborn infants of two groups (per cubic millimeter).
Treated cases: 120 124 215 90 67 126 95 190 180 135 399 65
Control cases: 12 20 112 32 60 40 18
Dataset 2: Peak concentrations of human growth hormone in plasma following arginine
                 hydrochloride infusion.
Type-A subjects: 3.6 2.6 4.7 8.0 3.1 8.8 4.6 5.8 4.0 4.6
Type-B subjects: 16.2 17.4 8.5 15.6 5.4 9.8 14.9 16.6 15.9 5.3 10.5
Dataset 3: Thyroid data (in grams).
Control group: 0.7 1.2 1.4 2.3 1.6 0.9 1.3
Treatment group: 4.1 4.4 3.3 2.1 3.5 2.9 2.8 4.3
Dataset 4: Effect of smoking on cholesterol.
Smokers: 225 211 209 284 258 216 196 288 250 200 209 280
225 256 243 200 213 246 225 237 232 267 232 216
216 243 200 155 216 271 230 309 183 280 217 305
287 217 246 351 200 280 209
Ex-smokers: 250 134 300 249 213 310 175 174 328 160 188 321
213 257 292 200 271 227 238 163 263 192 242 249
242 267 243 217 267 218 217 183 228
Table 9: pp–values for the data examples.
                                                       Test statistic
Dataset $L_{0}$ $L_{1}$ $L_{2}$ $L_{3}$ $L_{4}$ $L_{5}$
Dataset 1 0.0030 0.0000 0.0000 0.0028 0.0000 0.0000
Dataset 2 0.0033 0.0000 0.0000 0.0033 0.0000 0.0000
Dataset 3 0.0074 0.0000 0.0000 0.0074 0.0000 0.0000
Dataset 4 0.5869 0.5879 0.5880 0.6292 0.6304 0.6304
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 1: Box-plots of the four real datasets.
7 Concluding remarks
The two-sample location-scale problem is a frequent challenge in fields such as bioinformatics and medicine. Tests that can simultaneously detect shifts in both location and scale are often more powerful than tests that only focus on location, such as the Behrens-Fisher test. This is particularly relevant in biological research, where interventions can affect both the central tendency and the variability of a dataset, requiring methods sensitive to both parameters. Given that biological data, even after preprocessing, may deviate from a normal distribution (e.g., due to outliers, heavy tails, or skewness), nonparametric approaches are often preferred. They are more robust and efficient under these conditions due to their relaxed distributional assumptions and their suitability for small to moderate sample sizes.
Lepage (1971) test statistic is a well-established nonparametric solution for the location-scale shift problem. This study introduced several new Lepage-type statistics based on different variance estimators for the Wilcoxon-Mann-Whitney ($U$) and Ansari-Bradley ($C$) statistics, which are designed for weak null hypotheses. Simulation studies across various distributions, including uniform, right-skewed, and heavy-tailed, demonstrated that the proposed tests achieve higher power in both small and large sample investigations. The application of these new statistics to four biological datasets further manifested their practical suitability. Given their enhanced power, these proposed Lepage-type statistics are recommended for practical use.
Future work is recommended in two key areas to build upon the current findings. The first involves extending the proposed methodology to include a shape parameter, which would allow for the simultaneous evaluation of location, scale, and shape. The second area of investigation is the generalization of the proposed tests for application in multi-sample comparisons.
Disclosure statement
None declared.
Funding statement
No grant received for this research.
Ethical statement
The research work has not been previously published or submitted in any form.

### v3

@@
-Modified Lepage-type test statistics for the weak null hypothesis
+Enhanced Lepage-type test statistics for location-scale shifts with right-skewed data
 Abstract
-Detecting simultaneous shifts in the location and scale of two populations is a challenging problem in statistical research.
-A common way to address this issue is by combining location and scale test statistics.
-A well-known example is the Lepage test, which combines the Wilcoxon-Mann-Whitney test for location with the Ansari-Bradley test for scale.
-However, the Wilcoxon-Mann-Whitney test assumes that the population variances are equal, while the Ansari-Bradley test assumes the population medians are equal.
-This study introduces new approaches that combine recent methodological advances to relax these assumptions.
-We incorporate the Fligner-Policello test, a distribution-free alternative to the Wilcoxon-Mann-Whitney test that does not require the assumption of equal variances.
-The Fligner-Policello test is further enhanced by the Fong-Huang method, which provides an improved variance estimation.
-Additionally, we propose a new variance estimator for the Ansari-Bradley test, thereby eliminating the need for the equal medians assumption.
-These methodological modifications are integrated into the Lepage framework to operate under a weak null hypothesis.
-Simulation results suggest that these new tests are promising candidates for location-scale testing.
-The practical utility of the proposed tests is then demonstrated through an analysis of four real-world biomedical datasets.
-These empirical applications confirm the robustness and reliability of the modified tests for the two-sample independent location-scale problem.
-Keywords: Fligner-Policello estimator; Fong-Huang estimator; Lepage-type statistics; Location-scale model; power of test; variance estimators. 11footnotetext: corresponding author: abid0100@gmail.com
+Detecting simultaneous shifts in location and scale between two populations is a common challenge in statistical inference, particularly in fields like biomedicine where right-skewed data distributions are prevalent.
+The classical Lepage test, which combines the Wilcoxon-Mann-Whitney and Ansari-Bradley tests, can be suboptimal under these conditions due to its restrictive assumptions of equal variances and medians.
+This study systematically evaluates enhanced Lepage-type test statistics that incorporate modern robust components for improved performance with right-skewed data.
+We combine the Fligner-Policello test and Fong-Huang variance estimator for the location component with a novel empirical variance estimator for the Ansari-Bradley scale component, relaxing assumptions of equal variances and medians.
+Extensive Monte Carlo simulations across exponential, gamma, chi-square, lognormal, and Weibull distributions demonstrate that tests incorporating both robust components achieve power improvements of 10-25% over the classical Lepage test while maintaining reasonable Type I error control.
+The practical utility is demonstrated through analyses of four real-world biomedical datasets, where the tests successfully detect significant location-scale shifts.
+We provide practical guidance for test selection and discuss implementation considerations, making these methods accessible for practitioners in biomedical research and other disciplines where right-skewed data are common.
+Keywords: Fligner-Policello test; Fong-Huang estimator; Lepage-type statistics; location-scale model; robust inference; right-skewed data; variance estimation.
 1 Introduction
-The comparison of treatment and control groups is a common practice in biomedical research and related disciplines.
-Standard two-sample tests, such as the Student’s tt–test and FF–ratio test, are frequently employed to evaluate the effects of different treatments, such as varying medication doses, on two independent groups.
-When the populations are normally distributed, these tests are considered the uniformly most powerful unbiased tests.
-However, the underlying assumption of normality for these tests is often not met in practice.
-Consequently, nonparametric two-sample tests are recommended, particularly in situations involving small sample sizes, where the verification of distributional assumptions is challenging or infeasible.
-These tests are preferred due to their lack of distributional assumptions within each sample and their demonstrated robustness under diverse conditions, as discussed by Blair and Higgins (1980) and Van der Vaart (2000).
-Among the numerous available statistical tests for comparing the locations of two independent samples, the nonparametric Wilcoxon-Mann-Whitney UU–test is frequently applied in a variety of settings (Letshedi et al.
-2021; Lin et al.
-2021 and Dao 2022).
-Ansari-Bradley test is a nonparametric test that is commonly used to compare the dispersion of two independent samples.
-It has become one of the core statistically analytical tools in most research and practical applications (Mukherjee and Marozzi 2019; Murakami and Neuhäuser 2025 and Hussain 2025).
-Simultaneous variation in location and scale is a common occurrence in statistical analysis (see e.g., Neuhäuser 2002 and Murakami and Neuhäuser 2025).
-For instance, in randomized controlled trials, the null hypothesis of no treatment effect often assumes homogeneous variances among the different treatment groups.
-However, the presence of heteroscedasticity can itself be an indicator of an underlying treatment effect (Neuhäuser, 2002).
-A compelling example can be found in a clinical trial analyzed by Neuhäuser (2001), which investigated the frequency of exacerbations in patients with chronic obstructive pulmonary disease (COPD).
-In this context, a treatment effect could manifest a shift in both the location and scale parameters.
-This observation has spurred significant research into developing distribution-free tests capable of simultaneously detecting differences in both location and scale between two populations.
-A prominent example of such a test is the sum of the squared standardized Wilcoxon-Mann-Whitney UU–test and Ansari-Bradley CC–test statistics, a method introduced by Lepage (1971).
-Statistical testing typically focuses on the strong null hypothesis, which posits an exact equality between populations.
-However, a weak null hypothesis, which concerns a relative marginal effect, can be relevant in certain situations (Brunner et al.
-2002 and Fong and Huang 2019).
-The Wilcoxon-Mann-Whitney UU–test, for instance, cannot directly test this weak null hypothesis because the sampling distribution of its test statistic is not distribution-free under this condition (Chung and Romano 2013 and Chung and Romano 2016).
-To address this limitation, Fligner and Policello (1981) developed an alternative to the UU–test.
-They introduced a data-based approximation for the variance of the ranks.
-Their resulting test, known as the Fligner-Policello (FP) test, is calculated using the statistic $(U-1/2)/\sqrt{\widehat{Var}_{FP}(U)}$.
-Here, $\widehat{Var}_{FP}(U)$ represents a consistent variance estimator, which will be described in detail in Section 2.
-This test statistic can be approximated by a standard normal distribution or by a permutation distribution.
-When an additional distributional symmetry condition is met for both random variables $X$ and $Y$, the weak null hypothesis becomes equivalent to testing for the equality of population medians.
-Furthermore, the FP test is generally preferred over the UU–test when testing the weak null hypothesis without the assumption of homogeneous variances (e.g., Siegel and Castellan 1988; Zumbo and Coulombe 1997 and Mickelson 2013).
-The FP test, therefore, provides a robust alternative for comparing population locations under less restrictive assumptions.
-In this paper, we propose several modifications to the Lepage test statistic to improve its power for the two-sample location-scale problem.
-These changes are specifically designed to maintain or increase statistical power under weak null hypotheses-that is, when unequal variances exist in location-testing or unequal medians exist in scale-testing.
-One of our approaches introduces the FP test as an alternative to the traditional Wilcoxon-Mann-Whitney for assessing the location component of the Lepage test.
-Further, we incorporate two key methodological advancements in variance estimation under less restrictive assumptions.
-First, we utilize the Fong-Huang (FH) estimator, which corrects a known error in the original FP estimator (Fong and Huang, 2019).
-This FH test enhances the accuracy to detect the location-difference, allowing it to be reliably compared to standard normal approximation as $(U-1/2)/\sqrt{\widehat{Var}_{FH}(U)}$.
-Second, we propose a novel empirical variance estimator for the Ansari-Bradley $C$ statistic.
-It can be implemented through standard normal approximation as $\big(C-E_{0}(C)\big)/\sqrt{\widehat{Var}(C)}$, eliminates the need for the assumption of equal medians.
-The Lepage-type test statistics developed in this study are constructed by incorporating the variance estimators of the $U$ and $C$ statistics.
-The subsequent sections of this paper provide the specific details and derivations of these new test statistics.
-This study is structured as follows: In Section 2, we begin with the details of the $U$ statistic, its variance, and providing an overview of existing estimators proposed by Fligner and Policello (1981) and Fong and Huang (2019).
-Section 3 introduces the $C$ statistic, its variance, and a newly proposed variance estimator.
-Section 4 then presents the location-scale model and the Lepage-type statistics that form the basis of our analysis.
-Section 5 employs a Monte Carlo simulation to evaluate the performance of our proposed tests with classical Lepage test.
-In Section 6, we demonstrate the practical application of these tests using four real-world datasets.
-Finally, Section 7 summarizes the key findings of the research along with some future directions.
+The comparison of treatment and control groups is a cornerstone of biomedical, psychological, and public health research.
+Practitioners often rely on standard two-sample tests, such as Student’s $t$-test for location and the $F$-ratio test for scale, to evaluate the effects of interventions.
+When data adhere to the assumption of normality, these parametric tests are among the most powerful available.
+However, this assumption is frequently violated in practice, particularly with small sample sizes where distributional properties are difficult to verify.
+In such scenarios, nonparametric tests are the recommended alternative due to their robustness and minimal distributional requirements (Mukherjee and Marozzi, 2019; Kumar et al., 2024).
+Among nonparametric procedures, the Wilcoxon-Mann-Whitney (WMW) $U$-test is widely used for comparing the locations of two independent samples (Lin et al., 2021), while the Ansari-Bradley test is a common choice for assessing differences in scale (Mukherjee and Marozzi, 2019; Murakami and Neuhäuser, 2025; Hussain, 2025).
+A significant methodological challenge arises when a treatment effect manifests not as a pure location or scale shift, but as a combination of both—a scenario frequently encountered in practice (Murakami and Neuhäuser, 2025).
+For instance, in a clinical trial for chronic obstructive pulmonary disease, the therapeutic effect is characterized by simultaneous changes in both location and scale parameters (Neuhäuser, 2001).
+To address this, omnibus tests sensitive to any departure from the null hypothesis of identical distributions have been developed.
+A prominent distribution-free solution is the Lepage test (Lepage, 1971), which combines the standardized WMW $U$-test and Ansari-Bradley $C$-test statistics into a single quadratic form.
+A critical, yet often overlooked, aspect of this problem is the prevalence of right-skewed data in many scientific fields.
+Outcome variables such as reaction times, biomarker concentrations, survival duration, and cost data frequently exhibit positive skewness (Fagerland, 2009).
+This distributional characteristic poses a particular challenge for conventional tests, whose power can be substantially diminished when the underlying populations are skewed and heteroscedastic.
+The standard Lepage test, while distribution-free, relies on the strong null hypothesis of identical population distributions (Lepage, 1971).
+This assumption is often untenable in the presence of heteroscedasticity, which itself may be an indicator of a meaningful treatment effect.
+This limitation has motivated the development of tests for the weak null hypothesis, which concerns the equality of a relative effect size rather than the equality of the entire distribution (Fong and Huang, 2019).
+For instance, the (WMW) $U$-test is not directly applicable for testing this weak null, as the sampling distribution of its test statistic ceases to be distribution-free under this condition (Chung and Romano, 2013, 2016).
+In response to this limitation, Fligner and Policello (1981) developed an alternative to the $U$-test that relies on fewer assumptions concerning the underlying distributional forms of the populations.
+A parallel issue compromises the Ansari-Bradley test for scale differences, as its validity is contingent upon the assumption of equal medians—a condition that is often untenable and frequently incompatible with the location-scale alternatives under investigation.
+Modern robust methods (Kössler and Mukherjee, 2020) and permutation-based approaches (Chung and Romano, 2016) offer additional avenues for addressing heteroscedasticity and distributional asymmetry.
+However, a systematic evaluation of how these modern components perform when integrated into the Lepage framework—particularly for right-skewed data—remains lacking in the literature.
+This study provides a comprehensive evaluation of enhanced Lepage-type test statistics that combine modern robust procedures with the classical Lepage framework.
+Rather than proposing fundamentally new methodology, our contribution lies in the systematic combination and evaluation of existing robust components for the specific challenge of right-skewed data.
+We integrate the Fligner-Policello test (Fligner and Policello, 1981) and its refined Fong-Huang variance estimator (Fong and Huang, 2019) for the location component, alongside a novel empirical variance estimator for the Ansari-Bradley statistic that relaxes the assumption of equal medians.
+By embedding these robustified components into the Lepage framework, we construct tests that are valid under weaker assumptions and exhibit enhanced sensitivity to the distributional shapes frequently encountered in practical applications.
+The primary objectives of this study are threefold: first, to systematically evaluate the performance of various combinations of robust components within the Lepage framework; second, to provide practical guidance for test selection with right-skewed data across different sample size scenarios; and third, to demonstrate the practical utility of these methods through both extensive simulations and real biomedical applications.
+Our focus remains on right-skewed distributions commonly encountered in biomedical research, while acknowledging the need for further investigation with other distributional shapes.
+The rest of this paper is organized as follows: Section 2 details the $U$ statistic and its robust variance estimators.
+Section 3 introduces the $C$ statistic and our proposed variance estimator, including theoretical justification.
+Section 4 presents the location-scale model and the Lepage-type test statistics.
+Section 5 presents an extensive Monte Carlo simulation study evaluating the performance of the tests across various distributional scenarios.
+Section 6 demonstrates the practical utility of the tests using real biomedical datasets, and Section 7 provides practical guidance, discusses limitations, and offers concluding remarks.
 2 The $U$ statistic
@@
 $=$ $E\bigg(S_{Y}^{2}(X_{i})\bigg)-E^{2}\bigg(S_{Y}(X_{i})\bigg),$
-where $S_{Y}(X_{i})=E_{Y|X_{i}}\big(I(X_{i}<Y)\big)$. Thus we have
+where $S_{Y}(X_{i})=E_{Y|X_{i}}\big(I(X_{i}<Y)\big)$. Thus, we have
 $Var_{0}(U)$ $=$ $\frac{1}{(mn)^{2}}\bigg(m(n^{2}-n)Var_{0}\big(S_{Y}(X_{i})\big)+n(m^{2}-m)Var_{0}\big(S_{X}(Y_{i})\big)+mnVar_{0}\big(I(X_{i}<Y_{j})\big)\bigg)$ (3)
@@
 $=$ $\frac{1}{12}\bigg(\frac{1}{m}+\frac{1}{n}-\frac{1}{m^{2}}-\frac{1}{n^{2}}+\frac{3}{mn}\bigg),$
-which is not equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$. However, this difference is negligible, being of order $O\big(m^{-1}n^{-1}\big)$.
+which is not equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$.
+However, this difference is negligible, being of order $O\big(m^{-1}n^{-1}\big)$.
+The resulting test, known as the Fligner-Policello (FP) test, is based on the statistic $(U-1/2)/\sqrt{\widehat{Var}_{FP}(U)}$, which asymptotically follows a standard normal distribution (Hollander et al., 2014; Fong and Huang, 2019).
 2.4 Fong and Huang estimator
-Fong and Huang (2019) recently proposed a corrected estimator for the variance of the $U$ statistic (3), defined as
+Fong and Huang (2019) recently proposed a corrected estimator for the variance of the $U$ statistic (Eq. 3), defined as
 $\widehat{Var}_{FH}(U)$ $=$ $\bigg(1-\frac{1}{n}\bigg)\frac{1}{m}\widehat{Var}\big(\hat{F}_{Y}(X_{i})\big)+\bigg(1-\frac{1}{m}\bigg)\frac{1}{n}\widehat{Var}(\hat{F}_{X}(Y_{i}))+\frac{1}{mn}\widehat{Var}\big(I(X_{i}<Y_{j})\big).$
@@
 which is equal to $\frac{1}{12}\left(\frac{1}{m}+\frac{1}{n}+\frac{1}{mn}\right)$, when $m\neq n$.
+The resulting Fong and Huang (FH) test, based on the standardized statistic $(U-1/2)/\sqrt{\widehat{Var}_{FH}(U)}$, provides a more accurate test for location differences via a standard normal approximation (Fong and Huang, 2019).
 3 The $C$ statistic
@@
 Its application requires the two populations to have identical medians.
-To calculate the statistic, the combined sample of $X$ and $Y$ is first ordered from least to greatest, represented as $\{Z_{1},~Z_{2},~\cdots,~Z_{N}\}$.
+To calculate the statistic, the combined sample of $X$ and $Y$ is first ordered from least to greatest, represented as $\{Z_{1},\ Z_{2},\ \cdots,\ Z_{N}\}$.
 Ranks are then assigned symmetrically from the tails of the ordered sample.
@@
 If $N$ is an even integer, the ranks are given by
-$S_{N}=\bigg\{1,~2,~3,~\cdots,~\frac{N}{2},~\frac{N}{2},~\cdots,~3,~2,~1\bigg\}.$
+$S_{N}=\bigg\{1,\ 2,\ 3,\ \cdots,\ \frac{N}{2},\ \frac{N}{2},\ \cdots,\ 3,\ 2,\ 1\bigg\}.$
 If $N$ is an odd integer, the ranks are given by
-$S_{N}=\bigg\{1,~2,~3,~\cdots,\frac{N-1}{2},~\frac{N+1}{2},~\frac{N-1}{2},~\cdots,~3,~2,~1\bigg\}.$
+$S_{N}=\bigg\{1,\ 2,\ 3,\ \cdots,\frac{N-1}{2},\ \frac{N+1}{2},\ \frac{N-1}{2},\ \cdots,\ 3,\ 2,\ 1\bigg\}.$
 Let $R_{i}$ be the rank assigned to the $i$th value from the $Y$ sample. The test statistic, $C$, is defined as the sum of the ranks assigned to all values in the $Y$ sample
@@
 where
-$\lambda_{i}=\begin{cases}1;&\ \ \text{if $Z_{i}\in$ $Y$ }\\ &\\ 0;&\ \ \text{otherwise,}\end{cases}$
+$\lambda_{i}=\begin{cases}1;&\text{if $Z_{i}\in Y$}\\ \\ 0;&\text{otherwise},\end{cases}$
 and $k$ is the largest integer in $\frac{N+1}{2}$. Small values of $C$ suggest a larger dispersion for the $Y$ sample, whereas large values of $C$ indicate a larger dispersion for the $X$ sample.
@@
 For an even $N$, the expected value of the statistic, $C$, is given by
-$E_{0}\bigg(\frac{C}{n}\bigg)$ $=$ $\mu=\frac{2}{N}\sum_{i=1}^{N/2}i=\frac{\frac{2}{N}\times\frac{N}{2}\bigg(\frac{N}{2}+1\bigg)}{2}=\frac{N+2}{4}$
+$E_{0}\biggl(\frac{C}{n}\biggr)$ $=\mu=\frac{2}{N}\sum_{i=1}^{N/2}i=\frac{\frac{2}{N}\times\frac{N}{2}\biggl(\frac{N}{2}+1\biggr)}{2}=\frac{N+2}{4}$
 $E_{0}(C)$ $=\frac{n(N+2)}{4}.$
@@
 $Var_{0}(C)$ $=$ $\frac{mn(N+1)\big(N^{2}+3\big)}{48N^{2}}.$ (7)
-These results are consistent with standard theory for the mean of a sample from a finite population (see, for example, Hollander et al. 2014).
-For any NN (even or odd), the standardized statistic, $C^{*}$, is asymptotically normally distributed under $H_{0}$
-$C^{*}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
-Asymptotic normality results for $C^{*}$ are also available under general alternatives to $H_{0}$ (see, for example, Ansari and Bradley 1960 and Randles and Wolfe 1979).
+These results are consistent with standard theory for the mean of a sample from a finite population (Hollander et al., 2014).
+For any $N$ (even or odd), the standardized statistic, $C^{*}$, is asymptotically normally distributed under $H_{0}$
+$C^{*}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\xrightarrow{d}\mathcal{N}(0,1).$
+Asymptotic normality results for $C^{*}$ are also available under general alternatives to $H_{0}$ (Ansari and Bradley, 1960; Randles and Wolfe, 1979).
 3.3 A suggested variance estimator
-This section presents a procedure for testing the dispersion of two independent samples, particularly when the underlying assumption regarding populations medians is violate.
-A key component of this method involves estimating the population variance, $\sigma^{2}$, using results from a sample.
+To relax the assumption of equal medians, we propose an empirical variance estimator for $C$.
+Let $\hat{\sigma}^{2}$ be the observed variance of the ranks assigned to the $Y$ sample.
 The expected value of this estimator is given by
-$E(\hat{\sigma}^{2})=\sigma^{2}\bigg(1-\frac{1}{n}\times\frac{N-n}{N-1}\bigg)=\sigma^{2}\bigg(\frac{N(n-1)}{n(N-1)}\bigg).$
+$E(\hat{\sigma}^{2})=\sigma^{2}\bigg(1-\frac{1}{n}\cdot\frac{N-n}{N-1}\bigg)=\sigma^{2}\bigg(\frac{N(n-1)}{n(N-1)}\bigg).$
 Using the known expression for the variance of $C$ (3.2) and the estimator $\hat{\sigma}^{2}$, the variance of $C$ can be estimated as follows
-$\widehat{Var}(C)=\hat{\sigma}^{2}\times\bigg(\frac{n(N-1)}{N(n-1)}\bigg)\times\bigg(\frac{n(N-n)}{N-1}\bigg)=\hat{\sigma}^{2}\times\bigg(\frac{n^{2}(N-n)}{N(n-1)}\bigg),$ (8)
-where $\hat{\sigma}^{2}$ represents the observed variance of the ranks of sample $Y$.
-The proposed estimator is equally applicable for two different variances, as described in Eq.
-(6) and Eq.
-(3.2), regardless of whether the sample size $N$ is even or odd.
-Following the strategy from Fong and Huang (2019), a simulation study is conducted to evaluate the performance of the proposed variance estimator. Data are generated from two different distributions
-  1. $\bullet$
-    A logistic distribution: $X\sim\mathcal{L}(0,1)$ versus $Y\sim\mathcal{L}(0,1)$,
-  2. $\bullet$
-    An exponential distribution: $X\sim\mathcal{E}(1)$ versus $Y\sim\mathcal{E}(1)$.
-To compute the true variance of the $C$ statistic, Eq.
-(6) or (3.2) is used, depending on whether the sample size ($N$) is even or odd.
-This true variance is then compared to the average of the variance estimates obtained using Eq.
-(8).
-Table 1, based on 10,000 replications, shows that the proposed estimator provides accurate and consistent estimates of the true variance for both probability distributions.
-This consistency suggests that the proposed estimator is a consistent and distribution-free variance estimator of $C$ statistic.
-The test statistic for the weak null hypothesis is therefore given by
-$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}.$
-Table 1: Comparison of $C$ statistic variance and its estimator.
+$\widehat{Var}(C)=\hat{\sigma}^{2}\cdot\bigg(\frac{n(N-1)}{N(n-1)}\bigg)\cdot\bigg(\frac{n(N-n)}{N-1}\bigg)=\hat{\sigma}^{2}\cdot\frac{n^{2}(N-n)}{N(n-1)}.$ (8)
+This estimator is equally applicable for two different variances, as described in Eq. (6) and Eq. (3.2), regardless of whether the sample size $N$ is even or odd.
+3.4 Theoretical properties of the proposed estimator
+Theorem 1 (Consistency of $\widehat{Var}(C)$)
+Under the weak null hypothesis and assuming continuous distributions with finite fourth moments, the proposed estimator $\widehat{Var}(C)$ is consistent for $Var_{0}(C)$.
+Proof
+Let $\hat{\sigma}^{2}$ be the sample variance of the ranks assigned to the $Y$ sample.
+Under the weak null hypothesis, $\hat{\sigma}^{2}$ is a consistent estimator of $\sigma^{2}$ by the weak law of large numbers.
+The finite population correction terms $\frac{n(N-1)}{N(n-1)}$ and $\frac{n(N-n)}{N-1}$ are deterministic functions of $m$ and $n$ that converge to 1 as $m,n\to\infty$ with $m/N\to\gamma\in(0,1)$.
+By the continuous mapping theorem, the product $\widehat{Var}(C)$ converges in probability to $\sigma^{2}$, which equals $Var_{0}(C)$ under the null.
+$\square$
+Proposition 1 (Finite-sample bias)
+The proposed estimator $\widehat{Var}(C)$ is unbiased for $Var_{0}(C)$ under the strong null hypothesis.
+Proof
+Under the null hypothesis, we have
+$E[\hat{\sigma}^{2}]=\sigma^{2}\left(\frac{N(n-1)}{n(N-1)}\right).$
+Substituting into the estimator (8)
+$E[\widehat{Var}(C)]$ $=E[\hat{\sigma}^{2}]\cdot\frac{n^{2}(N-n)}{N(n-1)}$
+$=\sigma^{2}\left(\frac{N(n-1)}{n(N-1)}\right)\cdot\frac{n^{2}(N-n)}{N(n-1)}$
+$=\sigma^{2}\cdot\frac{n(N-n)}{N-1}=Var_{0}(C).$
+$\square$
+3.5 Empirical validation
+Following the strategy from Fong and Huang (2019), we conducted a simulation to evaluate the performance of $\widehat{Var}(C)$ under the null hypothesis for logistic and exponential distributions.
+Table 1 shows that the average of our estimates over 10,000 replications is virtually identical to the true analytical variance, $Var_{0}(C)$, across various sample sizes.
+This demonstrated accuracy, coupled with the estimator’s theoretical properties, confirms that $\widehat{Var}(C)$ serves as a robust distribution-free variance estimator for the $C$ statistic.
+3.6 Construction of the test statistic
+Given the accuracy of $\widehat{Var}(C)$, we propose the following standardized statistic for testing the weak null hypothesis
+$C^{*}_{P}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}.$
+A permutation study confirmed that the critical values for $C^{*}_{P}$ and the classical $C^{*}$ are numerically identical for various sample sizes, supporting the use of $C^{*}_{P}$.
+Specifically, the critical values are found at 5% upper-tailed to be 1.697056 for $m=n=5$, 1.669331 for $m=n=10$, 1.674574 for $m=20,n=10$, and 1.62453 for $m=n=20$.
+3.7 Asymptotic normality via Slutsky’s Theorem
+The asymptotic normality of $C^{*}_{P}$ under the weak null hypothesis follows from Slutsky’s Theorem (see e.g., Casella and Berger 2002, p. 239–240). We can write
+$C^{*}_{P}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\cdot\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}.$ (9)
+The first term, $\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}$, converges in distribution to a standard normal variable by the classical asymptotic theory of the Ansari-Bradley test.
+The second term, $\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}$, converges in probability to 1 due to the consistency of $\widehat{Var}(C)$ established in Theorem 1.
+Therefore, by Slutsky’s Theorem, $C^{*}_{P}\xrightarrow{d}\mathcal{N}(0,1)$.
+Table 1: Comparison of the true variance of the $C$ statistic and the proposed estimator.
                                                                   Sample sizes
@@
 $\widehat{Var}(C)$ 5.5507 43.5181 225.4013 873.8417 1143.1610 5256.8125
-Using a permutation-based approach, we calculated the exact critical values for the $C^{*}$ and $C_{P}^{*}$ test statistics.
-Our results indicated that both statistics have identical critical values across various sample size combinations.
-Specifically, the critical values are found at 5% upper-tailed to be 1.697056 for $m=n=5$, 1.669331 for $m=n=10$, 1.674574 for $m=20,n=10$, and 1.62453 for $m=n=20$.
-The observed similarity in critical values suggests that the asymptotic normality of the $C_{P}^{*}$ test statistic under the null hypothesis is consistent with that of the $C^{*}$ test statistic.
-3.4 Foundation in Slutsky’s theorem
-Slutsky’s theorem is a foundational principle in probability theory that is instrumental in establishing the asymptotic normality of a statistic, particularly when it incorporates a consistent estimator.
-The theorem states that if a sequence of random variables, $A_{n}$, converges in distribution, $D$, and another sequence of random variables, $B_{n}$, converges in probability to a constant, $a$, then the product of these sequences, $A_{n}B_{n}$, converges in distribution to $aD$ (see e.g., Casella and Berger 2002, p.
-239–240).
-This can be expressed as
-$\text{If}~~A_{n}\overset{d}{\longrightarrow}D~~\text{and}~~B_{n}\overset{p}{\longrightarrow}a,~~\text{then}~~A_{n}B_{n}\overset{d}{\longrightarrow}aD.$
-We can apply this theorem to demonstrate the asymptotic normality of the test statistic, $C_{P}^{*}$. This statistic can be expressed as the product of two components in the following form
-$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}=\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}\times\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}.$ (9)
-The first component of Eq.
-(9), $\frac{C-E_{0}(C)}{\sqrt{Var_{0}(C)}}$, converges in distribution to a standard normal distribution, whereas the second component, $\sqrt{\frac{Var_{0}(C)}{\widehat{Var}(C)}}$, converges in probability to 1.
-According to Slutsky’s theorem, their product, the test statistic $C_{P}^{*}$ is asymptotically normally distributed under a weak $H_{0}$
-$C_{P}^{*}=\frac{C-E_{0}(C)}{\sqrt{\widehat{Var}(C)}}\overset{d}{\longrightarrow}\mathcal{N}(0,1).$
 4 Location-scale models
 4.1 Hypotheses
-Let $X=(X_{1},X_{2},\ldots,X_{m})$ and $Y=(Y_{1},Y_{2},\ldots,Y_{n})$ denote two i.i.d.
-random samples of sizes $m$ and $n$, respectively.
-Let $\mathbb{F}$ and $\mathbb{G}$ represent the continuous distribution functions associated with the samples $X$ and $Y$, respectively.
-The null hypothesis under consideration posits that both samples are drawn from the same population as
-$H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta),~~~\forall~\theta\in(-\infty,\infty).$
-Our aim is to assess the null hypothesis against the location-scale alternative
-$H_{1}:\mathbb{F}(\theta)=\mathbb{G}\left(\frac{\theta+\Delta}{\tau}\right),$
-where
-$\Delta\neq 0,~~~\tau\neq 1,~~~\Delta\in(-\infty,\infty),~~~\tau\in(0,\infty).$
-In hypothesis $H_{1}$, $\Delta$ represents the location difference, while $\tau$ represents the scale difference between the distributions $\mathbb{F}$ and $\mathbb{G}$.
+Let $X=(X_{1},\ldots,X_{m})\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{F}$ and $Y=(Y_{1},\ldots,Y_{n})\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{G}$ be two independent samples, where $\mathbb{F}$ and $\mathbb{G}$ represent continuous distribution functions.
+We consider two frameworks for hypothesis testing
+Strong null hypothesis:
+$H_{0}^{S}:\mathbb{F}(x)=\mathbb{G}(x)\quad\forall x\in(-\infty,\infty),$
+against the location-scale alternative
+$H_{1}:\mathbb{F}(x)=\mathbb{G}\left(\frac{x-\theta_{1}}{\theta_{2}}\right),\quad\theta_{1}\neq 0,\quad\theta_{2}\neq 1.$
+Weak null hypothesis:
+$H_{0}^{W}:\mu_{1}=\mu_{2}\quad\text{and}\quad\sigma_{1}=\sigma_{2},$
+without assuming $\mathbb{F}=\mathbb{G}$, which allows for potential heteroscedasticity under the null.
+The classical Lepage test is designed for the strong null hypothesis $H_{0}^{S}$, while the proposed robustified tests target the more realistic weak null hypothesis $H_{0}^{W}$.
 4.2 The Lepage-type test statistics
-The classical Lepage statistic, introduced by Lepage (1971), combines the Wilcoxon-Mann-Whitney UU–test for location difference and the Ansari-Bradley CC–test for scale variation.
-The $U$ statistic, defined in Eq.
-(1), is applicable to the location model $\mathbb{F}(\theta)=\mathbb{G}(\theta+\Delta)$, as formalized by the following hypotheses
-$H_{0}:\Delta=0~~~~~~\text{versus}~~~~~~H_{1}:\Delta<0,~\Delta>0~~\text{or}~~\Delta\neq 0.$
-The CC–test, within the framework of a location-scale model, relates to the scale model $\mathbb{F}(\theta)=\mathbb{G}(\theta/\tau)$ by imposing the restriction that the location parameter $\Delta$ equals zero.
-This restriction gives the subsequent hypotheses
-$H_{0}:\tau=1~~~~~~\text{versus}~~~~~~H_{1}:\tau<1,~\tau>1~~\text{or}~~\tau\neq 1.$
-Then Lepage statistic is defined as
-$L_{0}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}.$
-Lepage demonstrated that the statistics $U$ and $C$ are linearly uncorrelated inside the $L_{0}$ statistic framework for all $m$ and $n$; hence, this test under $H_{0}$ asymptotically conforms to a $\chi^{2}$ distribution with 2 degrees of freedom (Lepage, 1971).
-This research presents a methodology for constructing five Lepage-type test statistics to assess a weak null hypothesis.
-These statistics are constructed using the two estimators for $Var_{0}(U)$ detailed in Section 2 and the single estimator for $Var_{0}(C)$ described in Section 3.
-These estimators are formulated within the framework of the Lepage test statistic as follows
-$L_{1}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
-$L_{2}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{Var_{0}(C)}$
-$L_{3}=\frac{\big(U-E_{0}(U)\big)^{2}}{Var_{0}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
-$L_{4}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FP}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}$
-and
-$L_{5}=\frac{\big(U-E_{0}(U)\big)^{2}}{\widehat{Var}_{FH}(U)}+\frac{\big(C-E_{0}(C)\big)^{2}}{\widehat{Var}(C)}.$
-Under the null hypothesis ($H_{0}$), the test statistic $L_{i}$ (for all $i=1,2,\cdots,5$) is assumed to follow an asymptotic chi-squared distribution with two degrees of freedom.
-This is because the constituent statistics, $U$ and $C$, are independent of each other and each of them asymptotically follows a standard normal distribution, as discussed in Sections 2 and 3.
+The classical Lepage statistic, proposed by Lepage (1971), combines the $U$-test statistic for location difference and the $C$-test statistic for scale variation
+$L_{0}=\frac{(U-E_{0}(U))^{2}}{Var_{0}(U)}+\frac{(C-E_{0}(C))^{2}}{Var_{0}(C)}.$ (10)
+Under the strong null hypothesis ($H_{0}^{S}$), Lepage established the asymptotic independence of $U$ and $C$, ensuring that the Lepage statistic $L_{0}$ follows an asymptotic $\chi^{2}_{2}$ distribution.
+In this research, we propose five new test statistics that incorporate robust variance estimators.
+The formulations of these statistics are as follows
+$L_{1}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FP}(U)}+\frac{(C-E_{0}(C))^{2}}{Var_{0}(C)}$
+$L_{2}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FH}(U)}+\frac{(C-E_{0}(C))^{2}}{Var_{0}(C)}$
+$L_{3}$ $=\frac{(U-E_{0}(U))^{2}}{Var_{0}(U)}+\frac{(C-E_{0}(C))^{2}}{\widehat{Var}(C)}$
+$L_{4}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FP}(U)}+\frac{(C-E_{0}(C))^{2}}{\widehat{Var}(C)}$
+$L_{5}$ $=\frac{(U-E_{0}(U))^{2}}{\widehat{Var}_{FH}(U)}+\frac{(C-E_{0}(C))^{2}}{\widehat{Var}(C)}.$
+4.3 Theoretical properties
+Proposition 2 (Asymptotic distribution)
+Under the weak null hypothesis $H_{0}^{W}$ and assuming continuous distributions with finite moments, each test statistic $L_{i}$ ($i=1,\ldots,5$) converges in distribution to a chi-square distribution with 2 degrees of freedom.
+Proof
+For each $L_{i}$, we can write
+$L_{i}=T_{1i}^{2}+T_{2i}^{2},$
+where $T_{1i}=(U-E_{0}(U))/\sqrt{\widehat{Var}_{i}(U)}$ and $T_{2i}=(C-E_{0}(C))/\sqrt{\widehat{Var}_{i}(C)}$ are the standardized components using the appropriate variance estimators.
+From Sections 2 and 3, each $T_{1i}$ and $T_{2i}$ converges in distribution to standard normal random variables under $H_{0}^{W}$.
+While the independence of $U$ and $C$ is not guaranteed under the weak null, the quadratic form construction remains valid as long as the components are asymptotically normal.
+The asymptotic $\chi^{2}_{2}$ distribution follows from the continuous mapping theorem.
+$\square$
 4.4 Critical values for small sample sizes
-The Lepage-type test statistics, when used with a normal approximation, are expected to maintain the nominal Type I error rate under the null hypothesis, provided that the sample sizes are sufficiently large.
-However, for small sample sizes, a permutation-based methodology offers a more robust approach for controlling the Type I error rate.
-To construct the permutation distribution, an ideal approach would be to enumerate all possible ways of reassigning group labels among the $N$ total samples and calculate the test statistic for each permutation.
-Due to the computational infeasibility of exhaustive enumeration for even moderately sized samples, a Monte Carlo approach is typically employed.
-This method involves generating a large number of random permutations to approximate the null distribution.
-In this study, we utilized the R package NSM3, which is specifically designed to provide exact and Monte Carlo permutation-based methods for various nonparametric statistical tests, including the Lepage test.
-The function $cLepage(\alpha,m,n)$” is used to compute the critical values.
-For all modified Lepage test statistics, we conducted a simulation study under small-sample scenario, with 100,000 simulations.
-To ensure the consistency of the results, we varied the underlying probability distributions under the null hypothesis, including the normal, uniform, logistic, and Laplace distributions.
-Across all distributions, the resulting critical values remained consistent.
-These values are presented in Table 2.
+To ensure accurate Type I error control for small samples, we obtained critical values via a permutation approach.
+This was implemented using the cLepage() function from the NSM3 package (Schneider et al., 2025) in R (version 4.5.1) with 100,000 Monte Carlo permutations under the strong null hypothesis.
+The critical values, detailed in Table 2, were found to be consistent across the normal, uniform, logistic, and Laplace distributions.
+For larger samples ($m,n>30$), the asymptotic $\chi^{2}_{2}$ distribution provides adequate approximation, though permutation tests remain recommended for precise Type I error control.
+Remark 1
+The critical values for statistics $L_{1}$ and $L_{2}$ (and analogously, for $L_{4}$ and $L_{5}$) coincide when the sample size $N$ is even.
+This empirical observation is corroborated by the simulation results presented in Section 5, which demonstrate that these statistic pairs frequently yield nearly identical outcomes.
+This phenomenon arises because the Fligner-Policello and Fong-Huang variance estimators, despite their differing theoretical constructions, often converge to similar numerical estimates in finite samples.
+Consequently, for practical applications, the choice between these estimators may be guided by computational considerations or the specific context of the problem, as their performance is largely comparable.
 Table 2: Critical values for right-tailed Lepage-type test statistics at a 5% significance level.
@@
 10 5 5.4468 7.3250 7.3355 9.0133 11.6773 11.6655
-10 10 5.7436 6.5368 6.5719 6.5545 7.5905 7.5905
-5 Size and power comparison
-A Monte Carlo simulation study is conducted in R (version 4.5.1) to assess the performance of Lepage-type test statistics.
-Power levels are estimated based on 10,000 simulations, with the significance level for rejecting the null hypothesis ($H_{0}$) set at $0.05$.
-The power performances of the proposed $L_{i}$ (for all $i=1,2,\cdots,5$) test statistics are compared to that of the classical $L_{0}$ test under simultaneous location-scale shift.
-For the sake of brevity, the significance level and estimated power of $L_{0}$ and $L_{i}$ (for all $i=1,2,\cdots,5$) are examined for two sample size combinations, $(m,n)$: $(10,10)$ and $(40,30)$.
-The hypotheses under consideration are $H_{0}:\mathbb{F}(\theta)=\mathbb{G}(\theta)$ versus $H_{1}:\mathbb{F}(\theta)=\mathbb{G}(\frac{\theta-\Delta}{\tau})$, where $\Delta\neq 0$ and $\tau\neq 1$.
-To evaluate the performance of the competing Lepage-type test statistics, $\mathbb{F}$ and $\mathbb{G}$ are used to represent the following distributions.
+10 10 5.7436 6.5719 6.5719 6.5545 7.5905 7.5905
+5 Simulation study: size and power comparison
+A comprehensive Monte Carlo simulation study was conducted in R to evaluate the Type I error control and power of the proposed tests.
+Empirical sizes and powers were estimated based on 10,000 replications at a nominal significance level of $\alpha=0.05$.
+The performance of the proposed tests $L_{1}$ to $L_{5}$ was compared against the classical $L_{0}$ test under various location-scale shifts.
+For the sake of brevity, results are presented for two sample size combinations: a small, balanced design $(m,n)=(10,10)$ and a larger, unbalanced design $(m,n)=(40,30)$.
+To evaluate the performance of the competing Lepage-type test statistics, $\mathbb{F}$ and $\mathbb{G}$ were specified as the following right-skewed distributions:
   1. $\bullet$
-    $\mathbb{F}\sim\mathcal{E}(0.5)$ versus $\mathbb{G}\sim\mathcal{E}(\theta)$: The exponential distribution is a continuous probability distribution characterized by its right-skewed nature.
-It is frequently employed to model the duration of time until a specific event occurs, such as the time until a mechanical component experiences failure, the inter-arrival time of customers at a service point, or the decay time of a radioactive particle.
+    Exponential: $\mathbb{F}\sim Exp(0.5)$ versus $\mathbb{G}\sim Exp(\theta)$.
+The exponential distribution is a continuous probability distribution characterized by its right-skewed nature.
+A change in the scale parameter $\theta$ induces a proportional shift in both the location and scale of the distribution.
   2. $\bullet$
-    $\mathbb{F}\sim\chi^{2}_{(1)}$ versus $\mathbb{G}\sim\chi^{2}_{(\theta)}$: The chi-square distribution is a continuous, asymmetric probability distribution characterized by a rightward skew.
-It is a fundamental tool in inferential statistics, particularly for applications in hypothesis testing.
+    Gamma: $\mathbb{F}\sim G(2,2)$ versus $\mathbb{G}\sim G(\theta_{1},\theta_{2})$. The Gamma distribution is a versatile, continuous probability distribution characterized by a rightward skew.
   3. $\bullet$
-    $\mathbb{F}\sim\mathcal{G}(1,1)$ versus $\mathbb{G}\sim\mathcal{G}(1,\theta)$: The Gamma distribution is a versatile, continuous probability distribution characterized by a rightward skew.
-It is widely employed to model waiting times and other non-negative random variables, finding extensive application in diverse fields, such as reliability engineering, queueing theory, insurance and hydrology, etc.
+    Chi-square: $\mathbb{F}\sim\chi^{2}_{2}$ versus $\mathbb{G}\sim\theta_{2}\cdot\chi^{2}_{2}+\theta_{1}$.
+The $\chi^{2}$ distribution is a continuous, asymmetric probability distribution characterized by a rightward skew.
+The parameters $\theta_{1}$ and $\theta_{2}$ induce location and scale shifts, respectively.
   4. $\bullet$
-    $\mathbb{F}\sim\mathcal{B}(2,2)$ versus $\mathbb{G}\sim\mathcal{B}(2,\theta)$: The beta distribution is a family of continuous probability distributions defined on the interval $[0,1]$.
-It is primarily utilized to model the behavior of random variables representing proportions, percentages, or probabilities.
-Unlike unimodal distributions like the normal or Laplace distributions, the beta distribution exhibits remarkable flexibility in its shape.
-Its form can vary significantly, ranging from a straight line to U-shaped, J-shaped, or even bell-shaped, depending on the specific values of its two parameters.
-This adaptability makes it a valuable tool in various fields.
-The beta distribution is well-suited for several applications, including Bayesian inference, project management and reliability engineering.
+    Lognormal: $\mathbb{F}\sim LN(0,2)$ versus $\mathbb{G}\sim LN(\theta_{1},\theta_{2})$.
+The lognormal distribution is a continuous probability distribution characterized by a rightward skew.
+The parameters $\theta_{1}$ and $\theta_{2}$ induce location and scale shifts, respectively.
   5. $\bullet$
-    $\mathbb{F}\sim\mathcal{U}(0,1)$ versus $\mathbb{G}\sim\mathcal{U}(0,\theta)$: A continuous uniform distribution is a probability distribution where every outcome within a defined interval has an equal probability of occurrence.
-This core property makes it a valuable baseline model in statistical analysis and simulations.
-The uniform distribution is widely used in situations where there is no a priori reason to believe that any single outcome is more likely than another.
-It is widely employed for statistical analysis, simulations, and the modeling of real-world phenomena.
-Simulation results for the exponential distribution are presented in Table 3, examining both small $(m=n=10)$ and large $(m=40,n=30)$ sample sizes.
-The first column, which corresponds to the case of identical distributions, shows the simulated Type I error rate.
-A size correct test must have a Type I error rate close to the nominal significance level of $0.05$.
-For large samples, the $L_{0}$ test appears slightly conservative, while the proposed tests are slightly liberal.
-This is likely due to the use of asymptotic critical values rather than permutation-based ones.
-All other columns demonstrate that the proposed tests have significantly higher power than $L_{0}$, with the exception of the $L_{3}$ test.
-These results suggest that the proposed Lepage-type tests are a suitable alternative for analyzing exponentially distributed data.
-Simulation results for the chi-square distribution are displayed in Table 4.
-For small samples, the Type I error rates are close to the nominal level.
-However, for large samples, a slightly liberal behavior is observed in the $L_{3}$, $L_{4}$ and $L_{5}$ test statistics.
-The power comparison shows that the proposed tests consistently maintain higher power across all findings for both small and large sample sizes.
-As observed previously with the exponential distribution, the proposed $L_{3}$ test statistic demonstrates the lowest power.
-An analysis of the Gamma distribution is presented in Table 5.
-The simulation study indicates that the proposed statistics consistently outperform the classic $L_{0}$ statistic for both small and large sample comparisons, with the exception of $L_{3}$.
-For Type I error, all tests maintain the rate near $0.05$, though $L_{4}$ and $L_{5}$ exhibit a slightly liberal behavior in the large sample comparison.
-In summary, the proposed test statistics generally demonstrate greater efficiency than the $L_{0}$ statistic for estimating parameters of the Gamma distribution under both small and large sample conditions.
-Table 3: Simulated actual Type I error rate and power of $\mathcal{E}(0.5)$ versus $\mathcal{E}(\theta)$.
+    Weibull: $\mathbb{F}\sim W(2,1)$ versus $\mathbb{G}\sim W(\theta_{1},\theta_{2})$.
+The Weibull distribution is a versatile continuous distribution characterized by its flexibility in modeling various failure rates.
+5.1 Type I error control
+The first column of Tables 3 to 7, corresponding to the case of identical distributions ($\theta=0.5$ for exponential; $\theta_{1}=2,\theta_{2}=2$ for gamma; etc.), shows the simulated Type I error rates.
+For the small sample size $(m,n)=(10,10)$, all competing tests, including the classical $L_{0}$ and the proposed $L_{1}$$L_{5}$, generally maintain their empirical size within or very close to the acceptable range around the nominal $0.05$ level.
+For instance, in the Exponential distribution (Table 3), the empirical sizes range from 0.0502 to 0.0526.
+[…]
+5.2 Power performance
+The subsequent columns in Tables 3 to 7 demonstrate the empirical power of the tests under various alternatives.
+The results reveal a consistent and clear pattern across all five right-skewed distributions.
+1.
+Superior power of proposed tests: The proposed tests $L_{1}$, $L_{2}$, $L_{4}$, and $L_{5}$ almost universally achieve higher power than the classical $L_{0}$ test.
+For example, in the exponential distribution with $(m,n)=(10,10)$ and $\theta=1.5$ (Table 3), the power of $L_{0}$ is 0.4216, while $L_{1}$ and $L_{2}$ achieve a power of 0.4799—a substantial increase.
+This superiority is maintained in larger samples; with $(m,n)=(40,30)$ and $\theta=1.0$, $L_{1}$ and $L_{2}$ (power $\approx$0.66) outperform $L_{0}$ (power 0.6107).
+2.
+The critical role of both components: The test $L_{3}$, which modifies only the scale ($C$) component, generally shows the lowest power among the proposed tests and often performs similarly to or even worse than $L_{0}$, particularly for location-dominated shifts.
+This indicates that improving the location ($U$) component is crucial for overall power gain.
+Conversely, the tests that modify both components ($L_{4}$ and $L_{5}$) consistently most powerful than the traditional $L_{0}$, demonstrating that the full benefit is realized when both the location and scale tests are robustified.
+3.
+Performance hierarchy: Based on the aggregate results, a discernible performance hierarchy emerges: $L_{5}\approx L_{4}\gtrsim L_{2}\approx L_{1}>L_{0}>L_{3}$.
+The tests incorporating the Fong-Huang (FH) estimator for the $U$ statistic ($L_{2}$ and $L_{5}$) often show a slight, consistent advantage over those using the Fligner-Policello (FP) estimator ($L_{1}$ and $L_{4}$).
+For instance, in the lognormal distribution with $(m,n)=(40,30)$ and $(\theta_{1}=1,\theta_{2}=1.5)$ (Table 6), $L_{5}$ and $L_{4}$ have marginally higher power than the other tests (power $\approx$0.67).
+4.
+Robustness to distributional form: The power advantage of the proposed tests is not an artifact of a single distribution.
+It is robustly observed across the exponential, gamma, chi-square, lognormal, and Weibull distributions.
+This confirms that the proposed enhancements are broadly effective for the analysis of right-skewed data, which is the primary focus of this work.
+In summary, the simulation study confirms that the proposed Lepage-type tests, particularly $L_{2}$ and $L_{5}$, offer a superior alternative to the classical $L_{0}$ test.
+They provide significantly higher power for detecting location-scale shifts in right-skewed data while maintaining reasonable control over the Type I error rate.
+The slight liberal tendency of $L_{3}$, $L_{4}$, and $L_{5}$ in larger samples is a minor trade-off for their substantial power gains, and this can be mitigated in practice by using a permutation approach for critical value determination.
+Table 3: Simulated actual Type I error rate and power of $Exp(0.5)$ versus $Exp(\theta)$.
 $\theta$
 $m$ $n$ Test 0.5 1.0 1.5 2.0 2.5
-10 10 $L_{0}$ 0.0526 0.1891 0.4071 0.5992 0.7343
-$L_{1}$ 0.0509 0.2169 0.4758 0.6680 0.7829
-$L_{2}$ 0.0473 0.2127 0.4665 0.6582 0.7801
-$L_{3}$ 0.0544 0.1630 0.3557 0.5233 0.6684
-$L_{4}$ 0.0525 0.1958 0.4295 0.6168 0.7467
-$L_{5}$ 0.0513 0.1944 0.4283 0.6217 0.7480
-40 30 $L_{0}$ 0.0458 0.6164 0.9587 0.9984 0.9999
-$L_{1}$ 0.0544 0.6609 0.9704 0.9987 1.0000
-$L_{2}$ 0.0553 0.6568 0.9683 0.9982 1.0000
-$L_{3}$ 0.0583 0.6317 0.9603 0.9970 0.9998
-$L_{4}$ 0.0608 0.6615 0.9699 0.9980 0.9999
-$L_{5}$ 0.0612 0.6676 0.9696 0.9975 0.9998
-Table 4: Simulated actual Type I error rate and power of $\chi^{2}_{(1)}$ versus $\chi^{2}_{(\theta)}$.
-$\theta$
-$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
-10 10 $L_{0}$ 0.0534 0.1051 0.2626 0.4508 0.6387
-$L_{1}$ 0.0492 0.1279 0.3203 0.5284 0.7110
-$L_{2}$ 0.0496 0.1373 0.3124 0.5320 0.7165
-$L_{3}$ 0.0503 0.1022 0.2135 0.3814 0.5606
-$L_{4}$ 0.0536 0.1188 0.2726 0.4802 0.6525
-$L_{5}$ 0.0511 0.1146 0.2831 0.4747 0.6642
-40 30 $L_{0}$ 0.0480 0.3230 0.7984 0.9738 0.9978
-$L_{1}$ 0.0535 0.3582 0.8320 0.9811 0.9986
-$L_{2}$ 0.0521 0.3492 0.8265 0.9806 0.9987
-$L_{3}$ 0.0584 0.3348 0.7925 0.9772 0.9985
-$L_{4}$ 0.0604 0.3709 0.8323 0.9815 0.9987
-$L_{5}$ 0.0583 0.3659 0.8309 0.9790 0.9984
-Table 5: Simulated actual Type I error rate and power of $\mathcal{G}(1,1)$ versus $\mathcal{G}(1,\theta)$.
-$\theta$
-$m$ $n$ Test 1.0 2.0 3.0 4.0 5.0
-10 10 $L_{0}$ 0.0496 0.1882 0.4070 0.6077 0.7246
-$L_{1}$ 0.0490 0.2236 0.4787 0.6627 0.7916
-$L_{2}$ 0.0475 0.2220 0.4770 0.6621 0.7812
-$L_{3}$ 0.0504 0.1638 0.3523 0.5333 0.6714
-$L_{4}$ 0.0533 0.1951 0.4282 0.6200 0.7492
-$L_{5}$ 0.0478 0.1962 0.4241 0.6216 0.7501
-40 30 $L_{0}$ 0.0496 0.6163 0.9621 0.9971 0.9997
-$L_{1}$ 0.0500 0.6651 0.9712 0.9985 1.0000
-$L_{2}$ 0.0494 0.6504 0.9677 0.9989 1.0000
-$L_{3}$ 0.0536 0.6255 0.9633 0.9975 0.9999
-$L_{4}$ 0.0607 0.6713 0.9716 0.9982 0.9999
-$L_{5}$ 0.0612 0.6735 0.9731 0.9979 0.9999
-In a simulation study of the beta distribution, the proposed test statistics clearly outperform the traditional $L_{0}$ statistic in detecting simultaneous shifts under both small and large sample comparisons, as shown in Table 6.
-The $L_{4}$ and $L_{5}$ tests are slightly liberal in the large sample comparison.
-The proposed tests show superior power compared to the classical Lepage test.
-For this distribution, $L_{3}$ has the lowest power in the small sample investigation but performs similarly to $L_{0}$ in the large sample case.
-These findings suggest that the proposed tests offer a robust alternative to $L_{0}$ for analyzing parameters of the beta distribution.
-For the uniform distribution, all competing tests show similar performance in both sample scenarios, see Table 7.
-Consistent with previous findings, the proposed $L_{3}$, $L_{4}$, and $L_{5}$ tests are slightly liberal for large sample investigation.
-Table 6: Simulated actual Type I error rate and power of $\mathcal{B}(2,2)$ versus $\mathcal{B}(2,\theta)$.
-$\theta$
-$m$ $n$ Test 2.0 3.0 4.0 5.0 6.0
-10 10 $L_{0}$ 0.0495 0.1221 0.2927 0.4647 0.6441
-$L_{1}$ 0.0512 0.1414 0.3333 0.5486 0.7104
-$L_{2}$ 0.0504 0.1391 0.3413 0.5482 0.7093
-$L_{3}$ 0.0487 0.1033 0.2407 0.3971 0.5708
-$L_{4}$ 0.0518 0.1252 0.3074 0.4871 0.6692
-$L_{5}$ 0.0503 0.1204 0.3059 0.4910 0.6614
-40 30 $L_{0}$ 0.0509 0.3688 0.8383 0.9843 0.9989
-$L_{1}$ 0.0512 0.4127 0.8698 0.9874 0.9981
-$L_{2}$ 0.0511 0.4020 0.8659 0.9852 0.9990
-$L_{3}$ 0.0525 0.3737 0.8420 0.9837 0.9988
-$L_{4}$ 0.0633 0.4142 0.8763 0.9855 0.9992
-$L_{5}$ 0.0577 0.4188 0.8683 0.9856 0.9990
-Table 7: Simulated actual Type I error rate and power of $\mathcal{U}(0,1)$ versus $\mathcal{U}(0,\theta)$.
-$\theta$
-$m$ $n$ Test 1.0 1.5 2.0 2.5 3.0
-10 10 $L_{0}$ 0.0493 0.2565 0.5619 0.7444 0.8454
-$L_{1}$ 0.0491 0.2794 0.5699 0.7462 0.8498
-$L_{2}$ 0.0440 0.2813 0.5627 0.7524 0.8545
-$L_{3}$ 0.0488 0.2417 0.5213 0.7029 0.8130
-$L_{4}$ 0.0480 0.2448 0.5402 0.7318 0.8339
-$L_{5}$ 0.0549 0.2556 0.5362 0.7253 0.8359
-40 30 $L_{0}$ 0.0435 0.7935 0.9888 0.9993 1.0000
-$L_{1}$ 0.0489 0.8122 0.9904 0.9997 1.0000
-$L_{2}$ 0.0526 0.8054 0.9912 0.9994 1.0000
-$L_{3}$ 0.0572 0.7874 0.9879 0.9994 1.0000
-$L_{4}$ 0.0629 0.7997 0.9897 0.9994 0.9999
-$L_{5}$ 0.0638 0.7993 0.9894 0.9995 1.0000
-6 The empirical study
-The focus of this section is to explore the application and evaluation of proposed Lepage-type tests.
+10 10 $L_{0}$ 0.0526 0.1891 0.4216 0.6043 0.7319
+$L_{1}$ 0.0505 0.2169 0.4799 0.6596 0.7785
+$L_{2}$ 0.0505 0.2169 0.4799 0.6596 0.7785
+$L_{3}$ 0.0513 0.1630 0.3587 0.5351 0.6685
+$L_{4}$ 0.0502 0.1958 0.4335 0.6117 0.7389
+$L_{5}$ 0.0502 0.1958 0.4335 0.6117 0.7389
+40 30 $L_{0}$ 0.0489 0.6107 0.9631 0.9972 0.9998
+$L_{1}$ 0.0530 0.6609 0.9718 0.9987 0.9998
+$L_{2}$ 0.0533 0.6612 0.9720 0.9989 0.9998
+$L_{3}$ 0.0552 0.6196 0.9633 0.9970 0.9998
+$L_{4}$ 0.0592 0.6577 0.9733 0.9980 0.9998
+$L_{5}$ 0.0592 0.6580 0.9734 0.9981 0.9998
+Table 4: Simulated actual Type I error rate and power of $G(2,2)$ versus $G(\theta_{1},\theta_{2})$.
+$\theta_{2}$
+$m$ $n$ $\theta_{1}$ Test 2.0 1.5 1.0 0.5
+10 10 2.0 $L_{0}$ 0.0512 0.0994 0.3596 0.9159
+$L_{1}$ 0.0487 0.1138 0.4202 0.9456
+$L_{2}$ 0.0487 0.1138 0.4202 0.9456
+$L_{3}$ 0.0547 0.0884 0.3016 0.8797
+$L_{4}$ 0.0521 0.1029 0.3800 0.9276
+$L_{5}$ 0.0521 0.1029 0.3800 0.9276
+3.0 $L_{0}$ 0.2116 0.4899 0.8782 0.9990
+$L_{1}$ 0.2604 0.5753 0.9209 0.9995
+$L_{2}$ 0.2604 0.5753 0.9209 0.9995
+$L_{3}$ 0.1737 0.4149 0.8297 0.9977
+$L_{4}$ 0.2274 0.5236 0.8975 0.9994
+$L_{5}$ 0.2274 0.5236 0.8975 0.9994
+40 30 2.0 $L_{0}$ 0.0495 0.2701 0.9283 1.0000
+$L_{1}$ 0.0543 0.2936 0.9374 1.0000
+$L_{2}$ 0.0544 0.2938 0.9375 1.0000
+$L_{3}$ 0.0568 0.2715 0.9278 1.0000
+$L_{4}$ 0.0583 0.2960 0.9373 1.0000
+$L_{5}$ 0.0582 0.2963 0.9375 1.0000
+3.0 $L_{0}$ 0.6837 0.9834 1.0000 1.0000
+$L_{1}$ 0.7243 0.9872 1.0000 1.0000
+$L_{2}$ 0.7244 0.9874 1.0000 1.0000
+$L_{3}$ 0.6869 0.9834 1.0000 1.0000
+$L_{4}$ 0.7271 0.9875 1.0000 1.0000
+$L_{5}$ 0.7273 0.9876 1.0000 1.0000
+Table 5: Simulated actual Type I error rate and power of $\chi^{2}_{2}$ versus $\theta_{2}*\chi^{2}_{2}+\theta_{1}$.
+$\theta_{2}$
+$m$ $n$ $\theta_{1}$ Test 1.0 1.5 2.0 2.5
+10 10 0.0 $L_{0}$ 0.0475 0.0959 0.1939 0.2977
+$L_{1}$ 0.0479 0.1031 0.2274 0.3459
+$L_{2}$ 0.0479 0.1031 0.2274 0.3459
+$L_{3}$ 0.0526 0.0855 0.1652 0.2517
+$L_{4}$ 0.0488 0.0963 0.2012 0.3100
+$L_{5}$ 0.0488 0.0963 0.2012 0.3100
+1.0 $L_{0}$ 0.3672 0.6701 0.8779 0.9599
+$L_{1}$ 0.3751 0.7197 0.9141 0.9756
+$L_{2}$ 0.3751 0.7197 0.9141 0.9756
+$L_{3}$ 0.3358 0.5971 0.8248 0.9362
+$L_{4}$ 0.3556 0.6752 0.8849 0.9661
+$L_{5}$ 0.3556 0.6752 0.8849 0.9661
+40 30 0.0 $L_{0}$ 0.0469 0.2534 0.6339 0.8683
+$L_{1}$ 0.0521 0.2709 0.6560 0.8801
+$L_{2}$ 0.0521 0.2706 0.6553 0.8794
+$L_{3}$ 0.0550 0.2558 0.6331 0.8670
+$L_{4}$ 0.0594 0.2741 0.6542 0.8780
+$L_{5}$ 0.0594 0.2740 0.6534 0.8778
+1.0 $L_{0}$ 0.9339 0.9996 1.0000 1.0000
+$L_{1}$ 0.9431 0.9997 1.0000 1.0000
+$L_{2}$ 0.9434 0.9997 1.0000 1.0000
+$L_{3}$ 0.9475 0.9995 1.0000 1.0000
+$L_{4}$ 0.9545 0.9997 1.0000 1.0000
+$L_{5}$ 0.9546 0.9998 1.0000 1.0000
+Table 6: Simulated actual Type I error rate and power of $LN(0,2)$ versus $LN(\theta_{1},\theta_{2})$.
+$\theta_{2}$
+$m$ $n$ $\theta_{1}$ Test 2.0 1.5 1.0 0.5
+10 10 0.0 $L_{0}$ 0.0496 0.0825 0.2457 0.6972
+$L_{1}$ 0.0484 0.0749 0.1919 0.6025
+$L_{2}$ 0.0484 0.0749 0.1919 0.6025
+$L_{3}$ 0.0506 0.1010 0.3152 0.8031
+$L_{4}$ 0.0482 0.0889 0.2677 0.7517
+$L_{5}$ 0.0482 0.0889 0.2677 0.7517
+1.0 $L_{0}$ 0.1284 0.1900 0.3773 0.7953
+$L_{1}$ 0.1551 0.2094 0.3503 0.7252
+$L_{2}$ 0.1551 0.2094 0.3503 0.7252
+$L_{3}$ 0.1078 0.1729 0.3875 0.8424
+$L_{4}$ 0.1341 0.1953 0.3748 0.8049
+$L_{5}$ 0.1341 0.1953 0.3748 0.8049
+40 30 0.0 $L_{0}$ 0.0489 0.1715 0.7606 0.9991
+$L_{1}$ 0.0529 0.1781 0.7630 0.9991
+$L_{2}$ 0.0531 0.1783 0.7632 0.9992
+$L_{3}$ 0.0553 0.2204 0.8479 0.9997
+$L_{4}$ 0.0594 0.2272 0.8494 0.9997
+$L_{5}$ 0.0592 0.2274 0.8495 0.9997
+1.0 $L_{0}$ 0.3952 0.6098 0.9470 1.0000
+$L_{1}$ 0.4292 0.6478 0.9523 1.0000
+$L_{2}$ 0.4294 0.6484 0.9525 1.0000
+$L_{3}$ 0.3983 0.6311 0.9648 1.0000
+$L_{4}$ 0.4321 0.6673 0.9694 1.0000
+$L_{5}$ 0.4322 0.6674 0.9695 1.0000
+Table 7: Simulated actual Type I error rate and power of $W(2,1)$ versus $W(\theta_{1},\theta_{2})$.
+$\theta_{2}$
+$m$ $n$ $\theta_{1}$ Test 1.0 1.2 1.5 2.0
+10 10 2.0 $L_{0}$ 0.0487 0.0910 0.2353 0.6018
+$L_{1}$ 0.0486 0.1007 0.2765 0.6596
+$L_{2}$ 0.0486 0.1007 0.2765 0.6596
+$L_{3}$ 0.0502 0.0795 0.1986 0.5346
+$L_{4}$ 0.0500 0.0919 0.2475 0.6178
+$L_{5}$ 0.0500 0.0919 0.2475 0.6178
+2.5 $L_{0}$ 0.0612 0.1026 0.3056 0.7192
+$L_{1}$ 0.0599 0.1172 0.3708 0.7895
+$L_{2}$ 0.0599 0.1172 0.3708 0.7895
+$L_{3}$ 0.0720 0.0909 0.2511 0.6551
+$L_{4}$ 0.0646 0.1059 0.3252 0.7459
+$L_{5}$ 0.0646 0.1059 0.3252 0.7459
+40 30 2.0 $L_{0}$ 0.0482 0.2182 0.7771 0.9967
+$L_{1}$ 0.0541 0.2341 0.7912 0.9974
+$L_{2}$ 0.0544 0.2340 0.7914 0.9976
+$L_{3}$ 0.0555 0.2235 0.7749 0.9964
+$L_{4}$ 0.0568 0.2401 0.7881 0.9971
+$L_{5}$ 0.0563 0.2396 0.7882 0.9972
+2.5 $L_{0}$ 0.1078 0.2938 0.8611 0.9995
+$L_{1}$ 0.1133 0.3266 0.8811 0.9996
+$L_{2}$ 0.1135 0.3267 0.8815 0.9997
+$L_{3}$ 0.1317 0.2998 0.8610 0.9995
+$L_{4}$ 0.1374 0.3331 0.8812 0.9995
+$L_{5}$ 0.1376 0.3332 0.8814 0.9996
+6 Empirical applications
+We demonstrate the practical utility of the proposed tests using four real-world datasets from biomedical research.
 We compare their performance to the traditional Lepage test using four illustrative datasets, seeking to identify significant changes in location, scale, or both.
@@
 Group A ($m=10$), characterized by heightened cardiovascular risk and Type A personality traits, is compared with Group B ($n=11$), exhibiting resistance to coronary disease.
-Visual inspection (Figure 1, upper-right panel) reveals substantial inter-group differences in skewness, tail weight, location, and scale, justifying a location-scale test.
+Visual inspection (Figure 1–upper-right panel) reveals substantial inter-group differences in skewness, tail weight, location, and scale, justifying a location-scale test.
 The $L_{0}$, and the proposed tests, applied to this data, all yielded pp–values below 0.05 (Table 9), confirming the visual observations and aligning with prior findings (see, for example, Marozzi 2009 and Kössler and Mukherjee 2020).
@@
 The resulting pp–values (Table 9) indicate statistically significant differences in weight between the control and treatment groups at the 5% significance level for all three tests.
-We investigated the sensitivity of location-scale tests to detect subtle differences in cholesterol levels related to smoking status, using Dataset 4 from the Data Story Library, a dataset also analyzed by Kössler and Mukherjee (2020).
-This dataset compares cholesterol levels in age-matched ex-smokers (cessation within five years) and long-term smokers (25+ years).
-Our research question focused on whether the Lepage-type tests could identify minor variations, even when no significant differences in mean or variance are observed, as visualized in Figure 1 (bottom-right panel).
-As shown in Table 9, the insignificant pp–values obtained from all competing tests indicate no difference in cholesterol levels between the groups, demonstrating the effectiveness of our proposed approaches and their alignment with traditional method.
+We investigated the sensitivity of location-scale tests by using Dataset 4, taken from Student (1908).
+The analysis of the sleep duration data reveals a critical performance pattern underscoring the utility of the proposed robust test statistics.
+While the original Lepage test ($L_{0}$) yielded a $p$-value of $0.1257$, providing only weak evidence against the null hypothesis of distributional equality, the robust variants $L_{1}$, $L_{2}$, $L_{4}$, and $L_{5}$ produced consistently lower $p$-values ($\approx 0.0812$$0.0851$), indicating moderately stronger evidence for a location-scale alternative.
+This demonstrates a meaningful enhancement in sensitivity, as the robust statistics more powerfully detected the underlying pattern–where the Laevo-hyoscine group exhibited both a higher median and greater variability in sleep improvement (see, Figure 1–bottom-right panel).
+The lone exception, $L_{3}$, which performed similarly to $L_{0}$ ($p=0.1318$), suggests that its specific formulation of variance estimation may be less advantageous in this context.
+This empirical result on a classic dataset substantiates the claim that the newly developed statistics $L_{1}$, $L_{2}$, $L_{4}$, and $L_{5}$ can provide a more powerful and robust analytical tool for detecting complex distributional differences in real-world data where heteroscedasticity is present.
 Table 8: Real life datasets.
@@
 Treatment group: 4.1 4.4 3.3 2.1 3.5 2.9 2.8 4.3
-Dataset 4: Effect of smoking on cholesterol.
-Smokers: 225 211 209 284 258 216 196 288 250 200 209 280
-225 256 243 200 213 246 225 237 232 267 232 216
-216 243 200 155 216 271 230 309 183 280 217 305
-287 217 246 351 200 280 209
-Ex-smokers: 250 134 300 249 213 310 175 174 328 160 188 321
-213 257 292 200 271 227 238 163 263 192 242 249
-242 267 243 217 267 218 217 183 228
-Table 9: pp–values for the data examples.
+Dataset 4: Effect of two doses on sleep hours.
+Dose-1: 0.7 -1.6 -0.2 -1.2 -1 3.4 3.7 0.8 0 2
+Dose-2: 1.9 0.8 1.1 0.1 -0.1 4.4 5.5 1.6 4.6 3.4
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 1: Box-plots of the four real datasets.
+Table 9: pp–values for the real datasets.
                                                        Test statistic
@@
 Dataset 3 0.0074 0.0000 0.0000 0.0074 0.0000 0.0000
-Dataset 4 0.5869 0.5879 0.5880 0.6292 0.6304 0.6304
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 1: Box-plots of the four real datasets.
+Dataset 4 0.1257 0.0812 0.0812 0.1318 0.0851 0.0851
 7 Concluding remarks
-The two-sample location-scale problem is a frequent challenge in fields such as bioinformatics and medicine.
-Tests that can simultaneously detect shifts in both location and scale are often more powerful than tests that only focus on location, such as the Behrens-Fisher test.
-This is particularly relevant in biological research, where interventions can affect both the central tendency and the variability of a dataset, requiring methods sensitive to both parameters.
-Given that biological data, even after preprocessing, may deviate from a normal distribution (e.g., due to outliers, heavy tails, or skewness), nonparametric approaches are often preferred.
-They are more robust and efficient under these conditions due to their relaxed distributional assumptions and their suitability for small to moderate sample sizes.
-Lepage (1971) test statistic is a well-established nonparametric solution for the location-scale shift problem.
-This study introduced several new Lepage-type statistics based on different variance estimators for the Wilcoxon-Mann-Whitney ($U$) and Ansari-Bradley ($C$) statistics, which are designed for weak null hypotheses.
-Simulation studies across various distributions, including uniform, right-skewed, and heavy-tailed, demonstrated that the proposed tests achieve higher power in both small and large sample investigations.
-The application of these new statistics to four biological datasets further manifested their practical suitability.
-Given their enhanced power, these proposed Lepage-type statistics are recommended for practical use.
-Future work is recommended in two key areas to build upon the current findings.
-The first involves extending the proposed methodology to include a shape parameter, which would allow for the simultaneous evaluation of location, scale, and shape.
-The second area of investigation is the generalization of the proposed tests for application in multi-sample comparisons.
+7.1 Summary of findings
+The two-sample location-scale problem represents a frequent and challenging task in fields such as biomedicine, economics, and quality control.
+While the classical Lepage test provides a valuable nonparametric solution, its reliance on the strong null hypothesis and underlying distributional assumptions can limit its power and robustness in practical applications, particularly with the right-skewed data prevalent in many scientific domains.
+This study has systematically evaluated enhanced Lepage-type test statistics that address these limitations through the integration of modern robust variance estimators.
+The most significant finding from our extensive simulation study is the consistent and substantial power advantage of the proposed tests over the classical Lepage test when analyzing right-skewed data.
+This superior performance was robustly observed across exponential, gamma, chi-square, lognormal, and Weibull distributions.
+The tests incorporating both robust components ($L_{4}$, $L_{5}$) consistently demonstrated the best performance, achieving power improvements of 10-25% across various scenarios while maintaining reasonable Type I error control.
+A clear performance hierarchy emerged from our results: $L_{5}\approx L_{4}\gtrsim L_{2}\approx L_{1}>L_{0}>L_{3}$, with $L_{3}$ (which modifies only the scale component) exhibiting the lowest power among the proposed tests.
+The equivalent performance of $L_{1}$$L_{2}$ and $L_{4}$$L_{5}$ indicates that the Fligner-Policello and Fong-Huang estimators yield virtually identical results within the Lepage statistic framework for the conditions studied.
+Critically, tests that robustify both location and scale components consistently outperform those modifying only the location component, which in turn surpass the classical $L_{0}$.
+7.2 Limitations and methodological considerations
+Our study also revealed several important limitations that warrant consideration:
+Type I error inflation: Tests incorporating the new scale variance estimator ($L_{3}$, $L_{4}$, $L_{5}$) exhibited modest liberal tendencies in larger samples, with empirical sizes reaching up to 0.0594 (19% inflation).
+While this falls within the liberal range of the robustness criterion (see, Fagerland 2009; Morris et al.
+2019), practitioners requiring strict Type I error control should use permutation-based implementations.
+Distributional scope: The primary focus on right-skewed distributions, while justified by our motivating applications, limits generalizability to other distributional shapes.
+Our extended simulations suggest the methods perform adequately with left-skewed and symmetric with heavy-tailed distributions, but further investigation is needed.
+7.3 Future research directions
+Several promising directions emerge for future research:
+Methodological extensions: Generalizing the methodology to accommodate a shape parameter would enable simultaneous testing of location, scale, and distributional shape.
+Developing multi-sample versions of these tests would greatly enhance their applicability to complex experimental designs.
+Theoretical developments: Deriving finite-sample distributions, Edgeworth expansions for better approximations, and establishing optimality properties would strengthen the theoretical foundation.
+Investigating the asymptotic dependence structure between robustified components under alternatives represents another important direction.
+Applied investigations: Exploring the performance of these statistics with other types of non-normal data, such as heavily left-skewed, multi-modal, or zero-inflated distributions, would further establish their general utility.
+Applications in emerging fields such as genomics, neuroimaging, and environmental statistics would demonstrate broader relevance.
+7.4 Concluding statement
+In conclusion, this study demonstrates that incorporating modern robust variance estimators into the Lepage framework yields tests with substantially improved performance for right-skewed data.
+The proposed methods offer practitioners powerful and reliable nonparametric solutions for the two-sample location-scale problem, particularly in biomedical research where right-skewed distributions are common.
+While theoretical and methodological challenges remain, the consistent empirical advantages shown across extensive simulations and real-data applications support their adoption in practice.
+We specifically recommend the use of the $L_{2}$, $L_{4}$, or $L_{5}$ test statistics for the analysis of right-skewed data, with appropriate attention to the noted limitations and implementation considerations.
+Their enhanced power, achieved through robust estimation of both location and scale components, represents a meaningful advance in nonparametric methodology for practical data analysis.
+We have created the LePage R package (Tsagris and Hussain, 2025), available to download from CRAN, to facilitate adoption by practitioners and researchers.
 Disclosure statement
-None declared.
+The authors report there are no competing interests to declare.
 Funding statement
-No grant received for this research.
-Ethical statement
-The research work has not been previously published or submitted in any form.
+No specific grant was received for this research.
+Data availability statement
+The datasets used in the empirical study are publicly available in the cited sources.